### v1

# Jacobian Descent For Multi-Objective Optimization

Abstract
Many optimization problems are inherently multi-objective. To address them, we formalize Jacobian descent (JD), a direct generalization of gradient descent for vector-valued functions. Each step of this algorithm relies on a Jacobian matrix consisting of one gradient per objective. The aggregator, responsible for reducing this matrix into an update vector, characterizes JD. While the multi-task learning literature already contains a variety of aggregators, they often lack some natural properties. In particular, the update should not conflict with any objective and should scale proportionally to the norm of each gradient. We propose a new aggregator specifically designed to satisfy this. Emphasizing conflict between objectives, we then highlight direct applications for our methods. Most notably, we introduce instance-wise risk minimization (IWRM), a learning paradigm in which the loss of each training example is considered a separate objective. On simple image classification tasks, IWRM exhibits promising results compared to the direct minimization of the average loss. The performance of our aggregator in those experiments also corroborates our theoretical findings. Lastly, as speed is the main limitation of JD, we provide a path towards a more efficient implementation.
[…]
1 Introduction
The field of multi-objective optimization studies minimization of vector-valued objective functions [1, 2, 3, 4]. In deep learning, a widespread approach to train a model with multiple objectives is to combine those into a scalar loss function minimized by stochastic gradient descent. While this method is simple, it comes at the expense of potentially degrading some individual objectives. Without prior knowledge of their relative importance, this is undesirable.
Early works have attempted to extend gradient descent (GD) to consider several objectives simultaneously, and thus several gradients [5, 6]. Essentially, [5, 6] propose some heuristic to prevent the degradation of any individual objective. Several other works have built upon this method, analyzing its convergence properties or extending it to a stochastic setting [7, 8, 9]. Later, this has been applied to multi-task learning to tackle conflict between tasks, illustrated by contradicting gradient directions [10]. Many studies have followed, proposing various other algorithms for the training of multi-task models [11, 12, 13, 14, 15, 16, 17]. They commonly rely on an aggregator that maps a collection of task-specific gradients (a Jacobian matrix) to a shared parameter update.
We propose to unify all such methods under the Jacobian descent (JD) algorithm, specified by an aggregator. [^1] Our library enabling JD with PyTorch is available at https://github.com/TorchJD/torchjd The goal of this algorithm is to minimize a differentiable vector-valued function ${\boldsymbol{f}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ iteratively without relying on a scalarization of the objective. Besides, we introduce a novel stochastic variant of JD that enables the training of neural networks with a large number of objectives. This unlocks a particularly interesting perspective: considering the minimization of instance-wise loss vectors rather than the usual minimization of the average training loss. As this paradigm is a direct generalization of the well-known empirical risk minimization (ERM) [18], we name it instance-wise risk minimization (IWRM).
With the JD formulation, we can ponder the properties that aggregators should satisfy. The need for a new aggregator emerges from the realization that none of the existing methods is both robust to conflict and proportionally affected by individual gradient magnitudes.
Our contributions are organized as follows: In Section 2, we formalize the JD algorithm and its stochastic variants. We then introduce three critical aggregator properties and define $\mathcal{A}_{\text{UPGrad}}$ to satisfy them. In the smooth convex case, we show convergence of JD with $\mathcal{A}_{\text{UPGrad}}$ to the Pareto front. We present applications for JD and aggregators in Section 3, emphasizing the IWRM paradigm. We then discuss existing aggregators and analyze their properties in Section 4. In Section 5, we report experiments with IWRM optimized with stochastic JD with various aggregators. Lastly, we address computational efficiency in Section 6, giving a path towards an efficient implementation.
2 Theoretical foundation
[…]
2.1 Jacobian descent
In the following, we introduce Jacobian descent, a natural extension of gradient descent supporting the optimization of vector-valued functions.
Suppose that ${\boldsymbol{f}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ is continuously differentiable. Let $\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\in\mathbb{R}^{m\times n}$ be the Jacobian matrix of ${\boldsymbol{f}}$ at ${\boldsymbol{x}}$, i.e. $\left[\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\right]_{ij}=\left[\frac{\partial}{\partial x_{j}}{\boldsymbol{f}}({\boldsymbol{x}})\right]_{i}$. If a given ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$ is updated with some ${\boldsymbol{y}}\in{\mathbb{R}}^{n}$, Taylor’s theorem yields
${\boldsymbol{f}}({\boldsymbol{x}}+{\boldsymbol{y}})={\boldsymbol{f}}({\boldsymbol{x}})+\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot{\boldsymbol{y}}+o(\|{\boldsymbol{y}}\|),$ (1)
where $o(\|{\boldsymbol{y}}\|)$ means that $\lim_{\|{\boldsymbol{y}}\|\to 0}\frac{{\boldsymbol{f}}({\boldsymbol{x}}+{\boldsymbol{y}})-{\boldsymbol{f}}({\boldsymbol{x}})-\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot{\boldsymbol{y}}}{\|{\boldsymbol{y}}\|}={\boldsymbol{0}}$. The term ${\boldsymbol{f}}({\boldsymbol{x}})+\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot{\boldsymbol{y}}$ is the first-order Taylor approximation of ${\boldsymbol{f}}({\boldsymbol{x}}+{\boldsymbol{y}})$. Since it depends on ${\boldsymbol{y}}$ only through $\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot{\boldsymbol{y}}$, it is sensible to make ${\boldsymbol{y}}$ a function of $\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})$. A mapping $\mathcal{A}:\mathbb{R}^{m\times n}\to{\mathbb{R}}^{n}$ reducing such a matrix into a vector is called an aggregator. For any $J\in\mathbb{R}^{m\times n}$, $\mathcal{A}(J)$ is called the aggregation of $J$ by $\mathcal{A}$. To minimize ${\boldsymbol{f}}$, the update is then selected as ${\boldsymbol{y}}=-\eta\mathcal{A}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}))$, where $\eta$ is an appropriate step size.
When $m=1$, the Jacobian consists of a single gradient. In GD, the update is thus simply $-\eta\nabla f({\boldsymbol{x}})$, i.e. the aggregator is the identity. A minimal version of GD is given in Algorithm 1. When $m>1$, the choice of $\mathcal{A}$ is non-trivial and is one of the main subjects of the present work. An elementary version of JD parameterized by some aggregator $\mathcal{A}$ is provided in Algorithm 2.
Input: ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$, $0<\eta$, $T\in\mathbb{N}$
for $t\leftarrow 1$ to $T$ do
       ${\boldsymbol{x}}\leftarrow{\boldsymbol{x}}-\eta\nabla f({\boldsymbol{x}})$
Output: ${\boldsymbol{x}}$
Algorithm 1 Gradient descent
Input: $\mathcal{A}:\mathbb{R}^{m\times n}\to{\mathbb{R}}^{n}$, ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$, $0<\eta$, $T\in\mathbb{N}$
for $t\leftarrow 1$ to $T$ do
       ${\boldsymbol{x}}\leftarrow{\boldsymbol{x}}-\eta\mathcal{A}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}))$
Output: ${\boldsymbol{x}}$
Algorithm 2 Jacobian descent with aggregator $\mathcal{A}$
Note that most gradient-based optimization algorithms, e.g. Adam [19], can similarly be extended to the multi-objective setting by substituting the gradient with an aggregation of the Jacobian.
In some settings, the exact computation of the update can be prohibitively slow or even intractable. When dealing with a single objective, the gradient $\nabla f({\boldsymbol{x}})$ can be substituted with an estimation. This is known as stochastic gradient descent (SGD). More generally, stochastic Jacobian descent (SJD) relies on estimates of the aggregation of the Jacobian. We highlight two straightforward methods for this. First, we can compute a stochastic estimation of the Jacobian and aggregate it instead of the true Jacobian. We call this stochastically estimated Jacobian descent (SEJD). Moreover, we can aggregate a matrix whose rows are a random subset of the rows of the true Jacobian. We refer to this as stochastic sub-Jacobian descent (SSJD). This novel approach enables multi-objective optimization with a very large number of objectives.
2.2 Desired properties for aggregators
An inherent challenge of multi-objective optimization is to manage conflicting objectives [10, 11, 12]. Note that substituting the update ${\boldsymbol{y}}=-\eta\mathcal{A}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}))$ into the first-order Taylor approximation ${\boldsymbol{f}}({\boldsymbol{x}})~{}+~{}\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot{\boldsymbol{y}}$ yields ${\boldsymbol{f}}({\boldsymbol{x}})-\eta\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot\mathcal{A}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}))$. In particular, if ${\boldsymbol{0}}\preceq\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot\mathcal{A}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}))$, then no coordinate of the approximation of ${\boldsymbol{f}}$ will increase. A pair of vectors ${\boldsymbol{x}},{\boldsymbol{y}}\in{\mathbb{R}}^{n}$ is said to conflict if ${\boldsymbol{x}}^{\top}{\boldsymbol{y}}<0$. Hence, if any row of $\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})$ conflicts with $\mathcal{A}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}))$, the corresponding coordinate of ${\boldsymbol{f}}$ will locally increase. When minimizing ${\boldsymbol{f}}$, avoiding conflict between the aggregation and any gradient is thus desirable. In GD, $\nabla{\boldsymbol{f}}({\boldsymbol{x}})$ does indeed not conflict with itself. For aggregators of JD, this motivates the first property.
Definition 1 (Non-conflicting).
Let $\mathcal{A}:\mathbb{R}^{m\times n}\to{\mathbb{R}}^{n}$ be an aggregator. If for all $J\in\mathbb{R}^{m\times n}$, ${\boldsymbol{0}}\preceq J\cdot\mathcal{A}(J)$, then $\mathcal{A}$ is said to be non-conflicting.
For any collection of vectors $C\subseteq{\mathbb{R}}^{n}$, the dual cone of $C$ is $\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}:\forall{\boldsymbol{y}}\in C,0\varleq{\boldsymbol{x}}^{\top}{\boldsymbol{y}}\}$ [20]. Notice that an aggregator $\mathcal{A}$ is non-conflicting if and only if for any $J$, $\mathcal{A}(J)$ is in the dual cone of the rows of $J$.
In a step of GD, the update scales proportionally to the gradient norm. Small gradients will thus lead to small updates, and conversely, large gradients will lead to large updates. Likewise, it would be coherent that the rows of the Jacobian also contribute to the aggregation proportionally to their norm. Scaling each row of a matrix $J\in\mathbb{R}^{m\times n}$ by the corresponding element of a vector ${\boldsymbol{c}}\in{\mathbb{R}}^{m}$ gives $\operatorname*{diag}({\boldsymbol{c}})\cdot J$. This insight can then be formalized as the following property.
Definition 2 (Linear under scaling).
Let $\mathcal{A}:\mathbb{R}^{m\times n}\to{\mathbb{R}}^{n}$ be an aggregator. If for all $J\in\mathbb{R}^{m\times n}$, the mapping from any ${\boldsymbol{0}}\prec{\boldsymbol{c}}\in{\mathbb{R}}^{m}$ to $\mathcal{A}(\operatorname*{diag}({\boldsymbol{c}})\cdot J)$ is linear in ${\boldsymbol{c}}$, then $\mathcal{A}$ is said to be linear under scaling.
Finally, (1) highlights that the precision of the first order Taylor approximation ${\boldsymbol{f}}({\boldsymbol{x}})+\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot{\boldsymbol{y}}$ improves as $\|{\boldsymbol{y}}\|$ decreases. For any candidate update ${\boldsymbol{y}}$, its projection ${\boldsymbol{y}}^{\prime}$ onto the row space of $\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})$ is such that $\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot{\boldsymbol{y}}^{\prime}=\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot{\boldsymbol{y}}$ and $\|{\boldsymbol{y}}^{\prime}\|\varleq\|{\boldsymbol{y}}\|$. This projection thus preserves the value of the approximation while improving its precision. It is therefore sensible to select ${\boldsymbol{y}}$ directly in the span of the rows of $\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})$, i.e. to have a vector of weights ${\boldsymbol{w}}\in{\mathbb{R}}^{m}$ such that ${\boldsymbol{y}}=\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})^{\top}\cdot{\boldsymbol{w}}$. This holds by design in GD and yields the last desired property for aggregators of JD.
Definition 3 (Weighted).
Let $\mathcal{A}:\mathbb{R}^{m\times n}\to{\mathbb{R}}^{n}$ be an aggregator. If for all $J\in\mathbb{R}^{m\times n}$, there exists ${\boldsymbol{w}}\in{\mathbb{R}}^{m}$ such that $\mathcal{A}(J)=J^{\top}\cdot{\boldsymbol{w}}$, then $\mathcal{A}$ is said to be weighted.
2.3 Unconflicting projection of gradients
We now define the unconflicting projection of gradients aggregator $\mathcal{A}_{\text{UPGrad}}$, specifically designed to be non-conflicting, linear under scaling and weighted. In essence, it projects each gradient onto the dual cone of the rows of the Jacobian and averages the results, as illustrated in Figure 1(a).
(a) $\mathcal{A}_{\text{UPGrad}}(J)$ (ours)
(b) $\mathcal{A}_{\text{Mean}}(J)$, $\mathcal{A}_{\text{MGDA}}(J)$ and $\mathcal{A}_{\text{DualProj}}(J)$
Figure 1: Aggregation of $J=[{\boldsymbol{g}}_{1}\;{\boldsymbol{g}}_{2}]^{\top}\in\mathbb{R}^{2\times 2}$ by four different aggregators. The dual cone of $\{{\boldsymbol{g}}_{1},{\boldsymbol{g}}_{2}\}$ is represented in green.
(a) $\mathcal{A}_{\text{UPGrad}}$ projects ${\boldsymbol{g}}_{1}$ and ${\boldsymbol{g}}_{2}$ onto the dual cone and averages the results.
(b) The mean $\mathcal{A}_{\text{Mean}}(J)=\frac{1}{2}({\boldsymbol{g}}_{1}+{\boldsymbol{g}}_{2})$ conflicts with ${\boldsymbol{g}}_{1}$. $\mathcal{A}_{\text{DualProj}}$ projects this mean onto the dual cone, so it lies on its frontier. $\mathcal{A}_{\text{MGDA}}(J)$ is less aligned with ${\boldsymbol{g}}_{2}$ because of its larger norm.
For any $J\in\mathbb{R}^{m\times n}$ and ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$, the projection of ${\boldsymbol{x}}$ onto the dual cone of the rows of $J$ is
$\pi_{J}({\boldsymbol{x}})=\operatorname*{arg\,min}_{\begin{subarray}{c}{\boldsymbol{y}}\in{\mathbb{R}}^{n}:~{}{\boldsymbol{0}}\preceq J{\boldsymbol{y}}\end{subarray}}\|{\boldsymbol{y}}-{\boldsymbol{x}}\|^{2}.$ (2)
Denoting by ${\boldsymbol{e}}_{i}\in{\mathbb{R}}^{m}$ the $i$th standard basis vector, $J^{\top}{\boldsymbol{e}}_{i}$ is the $i$th row of $J$. $\mathcal{A}_{\text{UPGrad}}$ is defined as
$\mathcal{A}_{\text{UPGrad}}(J)=\frac{1}{m}\sum_{i=1}^{m}\pi_{J}(J^{\top}{\boldsymbol{e}}_{i}).$ (3)
Since the dual cone is convex, any combination of its elements with positive coefficients remains in it. In particular, $\mathcal{A}_{\text{UPGrad}}(J)$ is always in the dual cone of the rows of $J$. $\mathcal{A}_{\text{UPGrad}}$ is thus non-conflicting. Note that if no pair of gradients conflicts, $\mathcal{A}_{\text{UPGrad}}$ simply averages the rows of the Jacobian.
Since $\pi_{J}$ is a projection onto a closed convex cone, if ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$ and $0<a\in\mathbb{R}$, then $\pi_{J}(a\cdot{\boldsymbol{x}})=a\cdot\pi_{J}({\boldsymbol{x}})$. By (3), $\mathcal{A}_{\text{UPGrad}}$ is thus linear under scaling.
When $n$ is large, the projection in (2) is prohibitively expensive to compute. An alternative but equivalent approach is to use its dual formulation, which is independent of $n$.
Proposition 1.
Let $J\in\mathbb{R}^{m\times n}$. For any ${\boldsymbol{u}}\in{\mathbb{R}}^{m}$, $\pi_{J}(J^{\top}{\boldsymbol{u}})=J^{\top}{\boldsymbol{w}}$ with
${\boldsymbol{w}}\in\operatorname*{arg\,min}_{\begin{subarray}{c}{\boldsymbol{v}}\in{\mathbb{R}}^{m}:~{}{\boldsymbol{u}}\preceq{\boldsymbol{v}}\end{subarray}}{\boldsymbol{v}}^{\top}JJ^{\top}{\boldsymbol{v}}.$ (4)
Proof.
See Appendix A.2. ∎
The problem defined in (4) can be solved efficiently using a quadratic programming solver, such as those bundled in qpsolvers [21]. For any $i\in[m]$, let ${\boldsymbol{w}}_{i}$ be given by (4) when substituting ${\boldsymbol{u}}$ with ${\boldsymbol{e}}_{i}$. Then, by Proposition 1,
$\mathcal{A}_{\text{UPGrad}}(J)=J^{\top}\left(\frac{1}{m}\sum_{i=1}^{m}{\boldsymbol{w}}_{i}\right).$ (5)
This provides an efficient implementation of $\mathcal{A}_{\text{UPGrad}}$ and proves that it is weighted. $\mathcal{A}_{\text{UPGrad}}$ can also be easily extended to incorporate a vector of preferences by replacing the averaging in (3) and (5) by a weighted sum with positive weights. This extension remains non-conflicting, linear under scaling and weighted.
2.4 Convergence
We now prove the convergence of JD with $\mathcal{A}_{\text{UPGrad}}$ when minimizing ${\boldsymbol{f}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ under standard assumptions. If for a given ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$, there exists no ${\boldsymbol{y}}\in{\mathbb{R}}^{n}$ such that ${\boldsymbol{f}}({\boldsymbol{y}})\precneq{\boldsymbol{f}}({\boldsymbol{x}})$, then ${\boldsymbol{x}}$ is said to be Pareto optimal. The set $X^{\ast}\subseteq{\mathbb{R}}^{n}$ of Pareto optimal points is called the Pareto set, and its image ${\boldsymbol{f}}(X^{\ast})$ is called the Pareto front.
Whenever ${\boldsymbol{f}}(\lambda{\boldsymbol{x}}+(1-\lambda){\boldsymbol{y}})\preceq\lambda{\boldsymbol{f}}({\boldsymbol{x}})+(1-\lambda){\boldsymbol{f}}({\boldsymbol{y}})$ holds for any pair of vectors ${\boldsymbol{x}},{\boldsymbol{y}}\in{\mathbb{R}}^{n}$ and any $\lambda\in[0,1]$, ${\boldsymbol{f}}$ is said to be -convex. Moreover, ${\boldsymbol{f}}$ is said to be $\beta$-smooth whenever $\|\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})-\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{y}})\|_{\text{F}}\ \varleq\beta\|{\boldsymbol{x}}-{\boldsymbol{y}}\|$ holds for any pair of vectors ${\boldsymbol{x}},{\boldsymbol{y}}\in{\mathbb{R}}^{n}$.
Theorem 1.
Let ${\boldsymbol{f}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ be a $\beta$-smooth and -convex function. Suppose that the Pareto front ${\boldsymbol{f}}(X^{\ast})$ is bounded and that for any ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$, there is ${\boldsymbol{x}}^{\ast}\in X^{\ast}$ such that ${\boldsymbol{f}}({\boldsymbol{x}}^{\ast})\preceq{\boldsymbol{f}}({\boldsymbol{x}})$. [^2] This condition is a generalization to the case $m\vargeq 1$ of the condition that there exists a minimizer ${\boldsymbol{x}}^{\ast}\in{\mathbb{R}}^{n}$. Let ${\boldsymbol{x}}_{0}\in{\mathbb{R}}^{n}$, and for all $t\in\mathbb{N}$, ${\boldsymbol{x}}_{t+1}={\boldsymbol{x}}_{t}-\eta\mathcal{A}_{\text{UPGrad}}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}_{t}))$, with $\eta=\frac{1}{\beta\sqrt{m}}$. Let ${\boldsymbol{w}}_{t}$ be the weights defining $\mathcal{A}_{\text{UPGrad}}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}_{t}))$ as per (5), i.e. $\mathcal{A}_{\text{UPGrad}}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}_{t}))=\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}_{t})^{\top}\cdot{\boldsymbol{w}}_{t}$. If ${\boldsymbol{w}}_{t}$ is bounded, then ${\boldsymbol{f}}({\boldsymbol{x}}_{t})$ converges to ${\boldsymbol{f}}({\boldsymbol{x}}^{\ast})$ for some ${\boldsymbol{x}}^{\ast}\in X^{\ast}$. In other words, ${\boldsymbol{f}}({\boldsymbol{x}}_{t})$ converges to the Pareto front.
Proof.
See Appendix A.3. ∎
Experimental considerations suggest that ${\boldsymbol{w}}_{t}$ converges to some ${\boldsymbol{w}}^{\ast}\in{\mathbb{R}}^{m}$ such that both ${\boldsymbol{0}}\prec{\boldsymbol{w}}^{\ast}$ and $\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}^{\ast})^{\top}{\boldsymbol{w}}^{\ast}={\boldsymbol{0}}$. This suggests that the boundedness of ${\boldsymbol{w}}_{t}$ could be relaxed or even removed from the set of assumptions of Theorem 1. This would demonstrate that JD with $\mathcal{A}_{\text{UPGrad}}$ and an appropriate step size converges to the Pareto front in the smooth convex case.
3 Applications
Instance-wise risk minimization
In machine learning, we generally have access to a training set consisting of $m$ examples. The goal of empirical risk minimization (ERM) [18] is simply to minimize the average loss over the whole training set. More generally, instance-wise risk minimization (IWRM) considers the loss associated with each training example as a distinct objective. Formally, if ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$ are the parameters of the model and $f_{i}({\boldsymbol{x}})$ is the loss associated to the $i$th example, the respective objective functions of ERM and IWRM are:
(Empirical risk) $\bar{f}({\boldsymbol{x}})$ $=\frac{1}{m}\sum_{i=1}^{m}f_{i}({\boldsymbol{x}})$ (6)
(Instance-wise risk) ${\boldsymbol{f}}({\boldsymbol{x}})$ $=\begin{bmatrix}f_{1}({\boldsymbol{x}})&f_{2}({\boldsymbol{x}})&\cdots&f_{m}({\boldsymbol{x}})\end{bmatrix}^{\top}$ (7)
Naively using GD for ERM is inefficient in most practical cases. A prevalent alternative is to use SGD or one of its variants. Similarly, using JD for IWRM is typically intractable. Indeed, it would require computing a Jacobian matrix with one row per training example at each iteration. In contrast, we can use the Jacobian of a random batch of training example losses. Since it consists of a subset of the rows of the full Jacobian, this approach is a form of stochastic sub-Jacobian descent, as introduced in Section 2.1. Note that IWRM can be extended to cases where each $f_{i}$ is a vector-valued function. The objective would then be the concatenation of the losses of all examples.
Multi-task learning
In multi-task learning, a single model is trained to perform several related tasks simultaneously, leveraging shared representations to improve overall performance [22]. At its core, multi-task learning is a multi-objective optimization problem [10], making it a straightforward application for Jacobian descent. Yet, conflict between tasks is often too limited to justify the overhead of computing all task-specific gradients, i.e. the whole Jacobian [23, 24]. In such cases, a practical approach is to minimize some linear scalarization of the objectives using an SGD-based method. Nevertheless, we believe that a setting with inherent conflict between tasks naturally prescribes Jacobian descent with a non-conflicting aggregator. We analyze several related works applied to multi-task learning in Section 4.
Adversarial training
In adversarial domain adaptation, the feature extractor of a model is trained with two conflicting objectives: The features should be helpful for the main task and should be unable to discriminate the domain of the input [25]. Likewise, in adversarial fairness, the feature extractor is trained to both minimize the predictability of sensitive attributes, such as race or gender, and maximize the performance on the main task [26]. Combining the corresponding gradients with a non-conflicting aggregator could enhance the optimization of such methods. We believe that the training of generative adversarial networks [27] could be similarly formulated as a multi-objective optimization problem. The generator and discriminator could then be jointly optimized with JD.
Momentum-based optimization
In gradient-based single-objective optimization, several methods use some form of gradient momentum to improve their convergence speed [28]. Essentially, their updates consider an exponential moving average of past gradients rather than just the last one. An appealing idea is to modify those algorithms to make them combine the gradient and the momentum with some aggregator, such as $\mathcal{A}_{\text{UPGrad}}$, rather than summing them. This would apply to many popular optimizers, like SGD with Nesterov momentum [29], Adam [19], AdamW [30] and NAdam [31].
Distributed optimization
In a distributed data-parallel setting with multiple machines or multiple GPUs, model updates are computed in parallel. This can be viewed as multi-objective optimization with one objective per data share. Rather than the typical averaging, a specialized aggregator, such as $\mathcal{A}_{\text{UPGrad}}$, could thus combine the model updates. This consideration can even be extended to federated learning, in which multiple entities participate in the training of a common model from their own private data by sharing model updates [32]. In this setting, as security is one of the main challenges, the non-conflicting property of the aggregator could be interesting or even necessary.
4 Existing aggregators
In the context of multi-task learning, several works have proposed iterative optimization algorithms based on the combination of task-specific gradients [10, 11, 13, 12, 14, 15, 16]. Aggregators can be extracted from these methods, thereby formulating them as variants of JD. More specifically, since the gradients are stochastically estimated from batches of data, these are cases of what we call SEJD. In the following, we briefly present the most prominent aggregators and summarize their properties in Table 1. We also consider $\mathcal{A}_{\text{Mean}}$, which simply averages the rows of the Jacobian. Their formal definitions are provided in Appendix B. Some of them are also illustrated in Figure 1(b).
$\mathcal{A}_{\text{RGW}}$ aggregates the matrix using a random vector of weights [14]. $\mathcal{A}_{\text{MGDA}}$ gives the aggregation that maximizes the smallest improvement [6, 10, 5]. $\mathcal{A}_{\text{CAGrad}}$ maximizes the smallest improvement in a ball around the average gradient whose radius is parameterized by $c\in\left[0,1\right[{}$ [12]. $\mathcal{A}_{\text{PCGrad}}$ projects each gradient onto the orthogonal hyperplane of other gradients in case of conflict, iteratively and in a random order [11]. It is, however, only non-conflicting when $m\varleq 2$, in which case $\mathcal{A}_{\text{PCGrad}}=m\cdot\mathcal{A}_{\text{UPGrad}}$. IMTL-G is a method to balance some gradients with impartiality [13]. It is only defined for linearly independent gradients, but we generalize it as a formal aggregator, denoted $\mathcal{A}_{\text{IMTL-G}}$, in Appendix B.6. Aligned-MTL orthonormalizes the Jacobian and weights its rows according to some preferences [16]. We denote by $\mathcal{A}_{\text{Aligned-MTL}}$ this method with uniform preferences. $\mathcal{A}_{\text{Nash-MTL}}$ aggregates Jacobians by finding the Nash equilibrium between task-specific gradients [15]. Lastly, the GradDrop layer [17] defines a custom backward pass that combines gradients with respect to some internal activation. The corresponding aggregator, denoted $\mathcal{A}_{\text{GradDrop}}$, randomly drops out some gradient coordinates based on their sign and sums the remaining ones.
In the context of continual learning, to limit forgetting, an idea is to project the gradient onto the dual cone of gradients computed with past examples [33]. This idea can be translated into an aggregator that projects the mean gradient onto the dual cone of the rows of the Jacobian. We name this $\mathcal{A}_{\text{DualProj}}$.
Several other works consider the gradients to be noisy when making their theoretical analysis [34, 35, 36, 37, 38]. Their solutions for combining gradients are typically stateful. Unifying their formulations would thus require a more complex variant of Jacobian descent.
In the federated learning setting, several aggregators have been proposed to combine the model updates while being robust to adversaries [39, 40, 41, 42]. We do not study them here as their desired properties mainly revolve around security.
Table 1: Properties satisfied for any number of objectives. Proofs are provided in Appendix B.
Ref. Aggregator Non- conflicting Linear under scaling Weighted
$\mathcal{A}_{\text{Mean}}$
[6, 10, 5] $\mathcal{A}_{\text{MGDA}}$
[33] $\mathcal{A}_{\text{DualProj}}$
[11] $\mathcal{A}_{\text{PCGrad}}$
[17] $\mathcal{A}_{\text{GradDrop}}$
[13] $\mathcal{A}_{\text{IMTL-G}}$
[12] $\mathcal{A}_{\text{CAGrad}}$
[14] $\mathcal{A}_{\text{RGW}}$
[15] $\mathcal{A}_{\text{Nash-MTL}}$
[16] $\mathcal{A}_{\text{Aligned-MTL}}$
(ours) $\mathcal{A}_{\text{UPGrad}}$
5 Experiments
In the following, we present empirical results for instance-wise risk minimization on some simple image classification datasets. IWRM is performed by stochastic sub-Jacobian descent, as described in Section 3. A key consideration is that when the aggregator is $\mathcal{A}_{\text{Mean}}$, this approach becomes equivalent to empirical risk minimization with SGD. Thus, it constitutes a baseline for comparison.
We train convolutional neural networks on subsets of SVHN [43], CIFAR-10 [44], EuroSAT [45], MNIST [46], Fashion-MNIST [47] and Kuzushiji-MNIST [48]. To make the comparisons as fair as possible, we have tuned the learning rate very precisely for each aggregator, as explained in detail in Appendix C.1. To keep computational costs reasonable, we have thus limited the size of each training dataset to 1024 images. This also enables re-running the same experiments several times on different subsets and with different seeds to gain confidence in our results. Note that this is strictly an optimization problem: we are not studying the generalization of the model, which would be captured by some performance metric on a test set. Other experimental settings, such as the network architectures and the total computational budget used to run our experiments, are given in Appendix C. Figure 2 reports the main results on SVHN and CIFAR-10, two of the datasets exhibiting the most substantial performance gap. Results on the other datasets and aggregators are reported in Appendix D.1. They also demonstrate a significant performance gap.
(a) SVHN: training loss
(b) SVHN: update similarity to the SGD update
(c) CIFAR-10: training loss
(d) CIFAR-10: update similarity to the SGD update
Figure 2: Optimization metrics obtained with IWRM with 1024 training examples and a batch size of 32, averaged over 8 random runs. The shaded area around each curve shows the estimated standard error of the mean over the 8 runs. Curves are smoothed for readability. Best viewed in color.
Here, we compare the aggregators in terms of their average loss over the training set. This is precisely the goal of ERM. For this reason, it is quite surprising to see that the direct optimization of this objective, as performed by $\mathcal{A}_{\text{Mean}}$, can be outperformed by some other aggregators. In particular, $\mathcal{A}_{\text{UPGrad}}$, and to some extent $\mathcal{A}_{\text{DualProj}}$, provide improvements on all datasets. Figures 2(b) and  2(d) show the similarity between the update of each aggregator and the update given by $\mathcal{A}_{\text{Mean}}$. As explained in Section 2.3, if there is no conflict among gradients, $\mathcal{A}_{\text{UPGrad}}$ becomes equivalent to $\mathcal{A}_{\text{Mean}}$. The similarity curve of $\mathcal{A}_{\text{UPGrad}}$ thus suggests substantial conflict between the individual gradients, especially early into the training. Our interpretation is that at the beginning, the gradients of hard examples are dominated by those of easier examples if averaged. However, with $\mathcal{A}_{\text{UPGrad}}$, the updates will positively affect hard examples even during the early training phase. Since fitting such examples is more complex and time-consuming, it is beneficial to consider them early on. Another possibility is that the projections of $\mathcal{A}_{\text{UPGrad}}$ improve its stability compared to $\mathcal{A}_{\text{Mean}}$, making it favor a learning rate giving larger updates, thus enabling a faster convergence. The sub-optimal performance of $\mathcal{A}_{\text{MGDA}}$ in this setting can be attributed to its sensitivity to small gradients. If any row of the Jacobian approaches zero, the aggregation by $\mathcal{A}_{\text{MGDA}}$ will also approach zero. Note that an advantage of linearity under scaling is to explicitly prevent this from happening.
Overall, these experiments demonstrate a high potential for the IWRM paradigm and confirm the validity of JD, and more specifically of SSJD, as multi-objective optimization algorithms. More generally, they question the soundness of the ubiquitous loss scalarization in deep learning, especially when conflict is significant. Besides, the superiority of $\mathcal{A}_{\text{UPGrad}}$ in such a simple setting confirms our theoretical findings on the desired properties of aggregators. Consequently, using $\mathcal{A}_{\text{UPGrad}}$ would likely be beneficial as an out-of-the-box replacement whenever some gradients are averaged or summed. An interesting consideration is that with SGD, increasing the batch size decreases the variance, but with SSJD combined with $\mathcal{A}_{\text{UPGrad}}$, the effect is non-trivial since it also narrows the dual cone. Additional results obtained when varying the batch size or when updating the parameters with the Adam optimizer are available in Appendices D.2 and D.3, respectively.
It is important to note that an iteration of SSJD takes more time than an iteration of SGD. The exact runtime of SSJD depends on many factors, such as the aggregator, the parallelization ability of the device on which the Jacobians are computed, and the implementation. We provide an empirical comparison of the speed of SGD and SSJD with different aggregators in Appendix E. Besides, we address computational efficiency concerns in the next section.
6 Towards an efficient implementation
When the number of objectives is dominated by the number of parameters of the model, the main overhead of JD comes from the computation of a Jacobian matrix rather than a single gradient. We give several paths towards faster algorithms.
Grouping objectives
Linear scalarization of a multi-objective optimization problem casts it as single-objective. This idea can be relaxed by considering $k$ different scalarizations of the objectives, with $k<m$ [49]. In the context of multi-task learning, several studies have already worked on task relationships and groupings [50, 51, 52, 53, 54]. We believe combining objectives with generally aligned gradients would be preferable.
Gramian-based JD
In the following, we show how to make a step of JD without even having to compute the Jacobian. For any $J\in\mathbb{R}^{m\times n}$, the matrix $G=JJ^{\top}$ is called the Gramian of $J$ and is positive semi-definite. Let $\mathcal{M}_{m}\subseteq\mathbb{R}^{m\times m}$ be the set of positive semi-definite matrices. The Gramian of the Jacobian, denoted $\mathcal{G}\kern-1.0pt{\boldsymbol{f}}({\boldsymbol{x}})=\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})^{\top}\in\mathcal{M}_{m}$, captures the relations – including conflicts – between all pairs of gradients.
Whenever $\mathcal{A}$ is a weighted aggregator, the update of JD is ${\boldsymbol{y}}=-\eta\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})^{\top}{\boldsymbol{w}}$ for some vector of weights ${\boldsymbol{w}}\in{\mathbb{R}}^{m}$. Substituting this into the Taylor approximation of (1) gives
${\boldsymbol{f}}({\boldsymbol{x}}+{\boldsymbol{y}})={\boldsymbol{f}}({\boldsymbol{x}})-\eta\mathcal{G}\kern-1.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot{\boldsymbol{w}}+o\left(\eta\sqrt{{\boldsymbol{w}}^{\top}\cdot\mathcal{G}\kern-1.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot{\boldsymbol{w}}}\right).$ (8)
This expression only depends on the Jacobian through its Gramian. It is thus sensible to focus on aggregators whose weights are only function of the Gramian. Denoting this function as $\mathcal{W}~{}:~{}\mathcal{M}_{m}\to{\mathbb{R}}^{m}$, those aggregators are such that $\mathcal{A}(J)=J^{\top}\cdot\mathcal{W}(G)$. Remarkably, all weighted aggregators of Table 1 can be expressed in this form. For such aggregators, substitution and linearity of differentiation [^3] For any ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$ and any ${\boldsymbol{w}}\in{\mathbb{R}}^{m}$, $\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})^{\top}{\boldsymbol{w}}=\nabla({\boldsymbol{w}}^{\top}{\boldsymbol{f}})({\boldsymbol{x}})$ then yield
$\mathcal{A}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}))=\nabla\Big{(}\mathcal{W}(\mathcal{G}\kern-1.0pt{\boldsymbol{f}}({\boldsymbol{x}}))^{\top}\cdot{\boldsymbol{f}}\Big{)}({\boldsymbol{x}}).$ (9)
After computing $\mathcal{W}(\mathcal{G}\kern-1.0pt{\boldsymbol{f}}({\boldsymbol{x}}))$, a step of JD would thus only require the backpropagation of a scalar function. The computational cost of applying $\mathcal{W}$ depends on the aggregator and is often dominated by the cost of computing the Gramian. We now outline an promising alternative algorithm for the latter.
Similarly to the backpropagation algorithm, the chain rule can be leveraged to propagate the Gramian. Let ${\boldsymbol{g}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{k}$ and ${\boldsymbol{f}}:{\mathbb{R}}^{k}\to{\mathbb{R}}^{m}$, then for any ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$,
$\mathcal{G}\kern-1.0pt({\boldsymbol{f}}\circ{\boldsymbol{g}})({\boldsymbol{x}})=\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{g}}({\boldsymbol{x}}))\cdot\mathcal{G}\kern-1.0pt{\boldsymbol{g}}({\boldsymbol{x}})\cdot\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{g}}({\boldsymbol{x}}))^{\top}.$ (10)
Besides, when the function has multiple inputs, the Gramian can be computed as a sum of individual Gramians. Let ${\boldsymbol{f}}:\mathbb{R}^{n_{1}+\cdots+n_{k}}\to{\mathbb{R}}^{m}$ and ${\boldsymbol{x}}=\begin{bmatrix}{\boldsymbol{x}}_{1}^{\top}&\cdots&{\boldsymbol{x}}_{k}^{\top}\end{bmatrix}^{\top}$. We can write $\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})$ as the concatenation of Jacobians $\begin{bmatrix}\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}({\boldsymbol{x}})&\cdots&\mathcal{J}\kern-2.0pt_{k}{\boldsymbol{f}}({\boldsymbol{x}})\end{bmatrix}$, where $\mathcal{J}\kern-2.0pt_{i}{\boldsymbol{f}}({\boldsymbol{x}})$ is the Jacobian of ${\boldsymbol{f}}$ with respect to ${\boldsymbol{x}}_{i}$ evaluated at ${\boldsymbol{x}}$. For any $i\in[m]$, let $\mathcal{G}\kern-1.0pt_{i}{\boldsymbol{f}}({\boldsymbol{x}})=\mathcal{J}\kern-2.0pt_{i}{\boldsymbol{f}}({\boldsymbol{x}})\mathcal{J}\kern-2.0pt_{i}{\boldsymbol{f}}({\boldsymbol{x}})^{\top}$. Then
$\mathcal{G}\kern-1.0pt{\boldsymbol{f}}({\boldsymbol{x}}_{1},\dots,{\boldsymbol{x}}_{k})=\sum_{i=1}^{k}\mathcal{G}\kern-1.0pt_{i}{\boldsymbol{f}}({\boldsymbol{x}}_{1},\dots,{\boldsymbol{x}}_{k}).$ (11)
When a function is made of compositions and concatenations of simple functions, those two rules enable efficient computation of the Gramian. In particular, let ${\boldsymbol{f}}:{\mathbb{R}}^{n}\times{\mathbb{R}}^{\ell}\to{\mathbb{R}}^{m}$ be some layer of a sequential neural network. Given an input ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$, some parameter ${\boldsymbol{p}}\in{\mathbb{R}}^{\ell}$, the output is given by ${\boldsymbol{y}}={\boldsymbol{f}}({\boldsymbol{x}},{\boldsymbol{p}})$. If $G\in\mathcal{M}_{n}$ is some Gramian associated to ${\boldsymbol{x}}$, then forwarding $G$ through ${\boldsymbol{f}}$ yields
$\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}({\boldsymbol{x}},{\boldsymbol{p}})\cdot G\cdot\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}({\boldsymbol{x}},{\boldsymbol{p}})^{\top}+\mathcal{G}\kern-1.0pt_{2}{\boldsymbol{f}}({\boldsymbol{x}},{\boldsymbol{p}}).$ (12)
Appendix F contains examples of Gramian forwarding for some usual layers. It is worth noting that the matrices of (12) can be very sparse and structured. For this reason, representing them as arrays can be inefficient. Finding the optimal representations in all practical cases and devising a full algorithm for an optimized computation of the Gramian remains an open challenge.
Still, these considerations have the potential to drastically reduce the computation time of JD and SSJD for some architectures. This would make our methods more extensively applicable and would thus be a significant step towards unlocking multi-objective optimization as a practical paradigm.
7 Conclusion
In this paper, we have defined the JD algorithm, characterized by an aggregator. Moreover, we have formulated desired aggregator properties and proposed $\mathcal{A}_{\text{UPGrad}}$ to alleviate the weaknesses of existing methods. Then, we have presented several possible applications for JD and aggregators. In particular, we have introduced IWRM, a novel learning paradigm minimizing the vector of per-instance losses. An empirical assessment of IWRM on some usual image classification datasets shows promising results and hints that it could be used as a general-purpose learning paradigm. As speed is the main limitation of JD, we have given a path towards an efficient implementation. We think that a fast algorithm to compute the Gramian of the Jacobian could unlock the full potential of JD. An important takeaway from our experiments is that considering an optimization problem in its entire dimensionality can be beneficial to its resolution. We thus hope that our work enables more multi-objective optimization problems to be considered as such.
Limitations and future directions
Our experimentation has some important limitations. First, we only evaluate JD on the minimization of the IWRM objective. It would be essential to develop proper benchmarks to compare aggregators on a wide variety of problems. Ideally, such problems should involve substantially conflicting objectives, e.g. multi-task learning with inherently competing or even adversarial tasks. Then, we have limited our scope to the comparison of optimization speeds, disregarding generalization. While this simplifies the experiments and makes the comparison precise and rigorous, optimization and generalization are sometimes intertwined. We thus believe that future works should focus on both aspects.
Acknowledgments
We would like to express our sincere thanks to Scott Pesme, Emre Telatar, Matthieu Buot de l’Épine, Adrien Vandenbroucque, Alix Jeannerot and Damian Dudzicz for their careful and thorough review. The many insightful discussions that we shared with them were essential to this project.
Appendix A Proofs
A.1 Supplementary theoretical results
Recall that a function ${\boldsymbol{f}}:{\mathbb{R}}^{n}\to\mathbb{R}^{m}$ is -convex if for all ${\boldsymbol{x}},{\boldsymbol{y}}\in\mathbb{R}^{n}$ and any $\lambda\in[0,1]$,
${\boldsymbol{f}}(\lambda{\boldsymbol{x}}+(1-\lambda){\boldsymbol{y}})\preceq\lambda{\boldsymbol{f}}({\boldsymbol{x}})+(1-\lambda){\boldsymbol{f}}({\boldsymbol{y}}).$
Lemma 1.
If ${\boldsymbol{f}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ is a continuously differentiable -convex function, then for any pair of vectors ${\boldsymbol{x}},{\boldsymbol{y}}\in{\mathbb{R}}^{n}$, $\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})({\boldsymbol{y}}-{\boldsymbol{x}})\preceq{\boldsymbol{f}}({\boldsymbol{y}})-{\boldsymbol{f}}({\boldsymbol{x}})$.
Proof.
$\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})({\boldsymbol{y}}-{\boldsymbol{x}})$ $=\lim_{\lambda\to 0^{+}}\frac{{\boldsymbol{f}}({\boldsymbol{x}}+\lambda({\boldsymbol{y}}-{\boldsymbol{x}}))-{\boldsymbol{f}}({\boldsymbol{x}})}{\lambda}$ $\left(\begin{subarray}{c}\text{differentiation}\end{subarray}\right)$
$\preceq\lim_{\lambda\to 0^{+}}\frac{{\boldsymbol{f}}({\boldsymbol{x}})+\lambda({\boldsymbol{f}}({\boldsymbol{y}})-{\boldsymbol{f}}({\boldsymbol{x}}))-{\boldsymbol{f}}({\boldsymbol{x}})}{\lambda}$ (-$\left(\begin{subarray}{c}\preceq\text{-convexity}\end{subarray}\right)$
$={\boldsymbol{f}}({\boldsymbol{y}})-{\boldsymbol{f}}({\boldsymbol{x}}),$
which concludes the proof. ∎
Lemma 2.
Let $J\in\mathbb{R}^{m\times n}$, let ${\boldsymbol{u}}\in\mathbb{R}^{m}$ and let ${\boldsymbol{x}}\in\mathbb{R}^{n}$, then
${\boldsymbol{u}}^{\top}J{\boldsymbol{x}}\varleq\|{\boldsymbol{u}}\|\cdot\|J\|_{\text{F}}\cdot\|{\boldsymbol{x}}\|$
Proof.
Let $J_{i}$ be the $i$th row of $J$, then
$({\boldsymbol{u}}^{\top}J{\boldsymbol{x}})^{2}$ $\varleq\|{\boldsymbol{u}}\|^{2}\cdot\|J{\boldsymbol{x}}\|^{2}$ $\left(\begin{subarray}{c}\text{Cauchy-Schwartz}\\ \text{inequality}\end{subarray}\right)$
$=\|{\boldsymbol{u}}\|^{2}\cdot\sum_{i=1}^{m}(J_{i}^{\top}{\boldsymbol{x}})^{2}$
$\varleq\|{\boldsymbol{u}}\|^{2}\cdot\sum_{i=1}^{m}\|J_{i}\|^{2}\cdot\|{\boldsymbol{x}}\|^{2}$ $\left(\begin{subarray}{c}\text{Cauchy-Schwartz}\\ \text{inequality}\end{subarray}\right)$
$=\|{\boldsymbol{u}}\|^{2}\cdot\|J\|_{\text{F}}^{2}\cdot\|{\boldsymbol{x}}\|^{2},$
which concludes the proof. ∎
Recall that a function ${\boldsymbol{f}}:{\mathbb{R}}^{n}\to\mathbb{R}^{m}$ is $\beta$-smooth if for all ${\boldsymbol{x}},{\boldsymbol{y}}\in\mathbb{R}^{n}$,
$\|\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})-\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{y}})\|_{\text{F}}\varleq\beta\|{\boldsymbol{x}}-{\boldsymbol{y}}\|$ (13)
Lemma 3.
Let ${\boldsymbol{f}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ be $\beta$-smooth, then for any ${\boldsymbol{w}}\in{\mathbb{R}}^{m}$ and any ${\boldsymbol{x}},{\boldsymbol{y}}\in{\mathbb{R}}^{n}$,
$)\displaystyle{\boldsymbol{w}}^{\top}({\boldsymbol{f}}({\boldsymbol{x}})-{\boldsymbol{f}}({\boldsymbol{y}})-\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{y}})({\boldsymbol{x}}-{\boldsymbol{y}})$$)\varleq\frac{\beta}{2}\|{\boldsymbol{w}}\|\cdot\|{\boldsymbol{x}}-{\boldsymbol{y}}\|^{2}$ (14)
Proof.
${\boldsymbol{w}}^{\top}({\boldsymbol{f}}({\boldsymbol{x}})-{\boldsymbol{f}}({\boldsymbol{y}})-\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{y}})({\boldsymbol{x}}-{\boldsymbol{y}}))$
[…]
which concludes the proof. ∎
A.2 Proposition 1
Proposition 1.
Let $J\in\mathbb{R}^{m\times n}$. For any ${\boldsymbol{u}}\in{\mathbb{R}}^{m}$, $\pi_{J}(J^{\top}{\boldsymbol{u}})=J^{\top}{\boldsymbol{w}}$ with
${\boldsymbol{w}}$ $\in\operatorname*{arg\,min}_{\begin{subarray}{c}{\boldsymbol{v}}\in{\mathbb{R}}^{m}:~{}{\boldsymbol{u}}\preceq{\boldsymbol{v}}\end{subarray}}{\boldsymbol{v}}^{\top}JJ^{\top}{\boldsymbol{v}}.$ (4)
Proof.
This is a direct consequence of Lemma 4. ∎
Lemma 4.
Let $J\in\mathbb{R}^{m\times n}$, $G=JJ^{\top}$, ${\boldsymbol{u}}\in{\mathbb{R}}^{m}$. For any ${\boldsymbol{w}}\in{\mathbb{R}}^{m}$ satisfying
${\boldsymbol{u}}\preceq{\boldsymbol{w}}$ (15a)
${\boldsymbol{0}}\preceq G{\boldsymbol{w}}$ (15b)
${\boldsymbol{u}}^{\top}G{\boldsymbol{w}}={\boldsymbol{w}}^{\top}G{\boldsymbol{w}}$ (15c)
we have $\pi_{J}(J^{\top}{\boldsymbol{u}})=J^{\top}{\boldsymbol{w}}$. Such a ${\boldsymbol{w}}$ is the solution to
${\boldsymbol{w}}$ $\in\operatorname*{arg\,min}_{{\boldsymbol{u}}\preceq{\boldsymbol{v}}}{\boldsymbol{v}}^{\top}G{\boldsymbol{v}}.$
Proof.
Note that
$\pi_{J}(J^{\top}{\boldsymbol{u}})=\operatorname*{arg\,min}_{\begin{subarray}{c}{\boldsymbol{x}}\in{\mathbb{R}}^{n}:\\ {\boldsymbol{0}}\preceq J{\boldsymbol{x}}\end{subarray}}\frac{1}{2}\|{\boldsymbol{x}}-J^{\top}{\boldsymbol{u}}\|^{2}$
is a convex program. Consequently, the KKT conditions are both necessary and sufficient. The Lagragian is given by $\mathcal{L}({\boldsymbol{x}},{\boldsymbol{v}})=\frac{1}{2}\|{\boldsymbol{x}}-J^{\top}{\boldsymbol{u}}\|^{2}-{\boldsymbol{v}}^{\top}J{\boldsymbol{x}}$. The KKT conditions are then given by
$\begin{cases}\nabla_{\boldsymbol{x}}\mathcal{L}({\boldsymbol{x}},{\boldsymbol{v}})={\boldsymbol{0}}\\ {\boldsymbol{0}}\preceq{\boldsymbol{v}}\\ {\boldsymbol{0}}\preceq J{\boldsymbol{x}}\\ 0={\boldsymbol{v}}^{\top}J{\boldsymbol{x}}\end{cases}$
[…]
The simple change of variable ${\boldsymbol{w}}={\boldsymbol{u}}+{\boldsymbol{v}}$ finishes the proof of the first part.
Since ${\boldsymbol{x}}=J^{\top}({\boldsymbol{u}}+{\boldsymbol{v}})$, the Wolfe dual program of $\pi_{J}(J^{\top}{\boldsymbol{u}})$ gives
${\boldsymbol{w}}$ $\in{\boldsymbol{u}}+\operatorname*{arg\,max}_{{\boldsymbol{v}}\in{\mathbb{R}}^{m}:~{}{\boldsymbol{0}}\preceq{\boldsymbol{v}}}\mathcal{L}(J^{\top}({\boldsymbol{u}}+{\boldsymbol{v}}),{\boldsymbol{v}})$
$={\boldsymbol{u}}+\operatorname*{arg\,max}_{{\boldsymbol{v}}\in{\mathbb{R}}^{m}:~{}{\boldsymbol{0}}\preceq{\boldsymbol{v}}}\frac{1}{2}\|J^{\top}{\boldsymbol{v}}\|^{2}-{\boldsymbol{v}}^{\top}JJ^{\top}({\boldsymbol{u}}+{\boldsymbol{v}})$
$={\boldsymbol{u}}+\operatorname*{arg\,max}_{{\boldsymbol{v}}\in{\mathbb{R}}^{m}:~{}{\boldsymbol{0}}\preceq{\boldsymbol{v}}}-\frac{1}{2}{\boldsymbol{v}}^{\top}G{\boldsymbol{v}}-{\boldsymbol{v}}^{\top}G{\boldsymbol{u}}$
$={\boldsymbol{u}}+\operatorname*{arg\,min}_{{\boldsymbol{v}}\in{\mathbb{R}}^{m}:~{}{\boldsymbol{u}}\preceq{\boldsymbol{u}}+{\boldsymbol{v}}}\frac{1}{2}({\boldsymbol{u}}+{\boldsymbol{v}})^{\top}G({\boldsymbol{u}}+{\boldsymbol{v}})$
$=\operatorname*{arg\,min}_{{\boldsymbol{v}}^{\prime}\in{\mathbb{R}}^{m}:~{}{\boldsymbol{u}}\preceq{\boldsymbol{v}}^{\prime}}\frac{1}{2}{\boldsymbol{v}}^{\prime\top}G{\boldsymbol{v}}^{\prime},$
which concludes the proof. ∎
A.3 Theorem 1
Theorem 1.
Let ${\boldsymbol{f}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ be a $\beta$-smooth and -convex function. Suppose that the Pareto front ${\boldsymbol{f}}(X^{\ast})$ is bounded and that for any ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$, there is ${\boldsymbol{x}}^{\ast}\in X^{\ast}$ such that ${\boldsymbol{f}}({\boldsymbol{x}}^{\ast})\preceq{\boldsymbol{f}}({\boldsymbol{x}})$. Let ${\boldsymbol{x}}_{0}\in{\mathbb{R}}^{n}$, and for all $t\in\mathbb{N}$, ${\boldsymbol{x}}_{t+1}={\boldsymbol{x}}_{t}-\eta\mathcal{A}_{\text{UPGrad}}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}_{t}))$, with $\eta=\frac{1}{\beta\sqrt{m}}$. Let ${\boldsymbol{w}}_{t}$ be the weights defining $\mathcal{A}_{\text{UPGrad}}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}_{t}))$ as per (5), i.e. $\mathcal{A}_{\text{UPGrad}}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}_{t}))=\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}_{t})^{\top}\cdot{\boldsymbol{w}}_{t}$. If ${\boldsymbol{w}}_{t}$ is bounded, then ${\boldsymbol{f}}({\boldsymbol{x}}_{t})$ converges to ${\boldsymbol{f}}({\boldsymbol{x}}^{\ast})$ for some ${\boldsymbol{x}}^{\ast}\in X^{\ast}$. In other words, ${\boldsymbol{f}}({\boldsymbol{x}}_{t})$ converges to the Pareto front.
To prove the theorem we will need Lemmas 5, 6 and 7 below.
Lemma 5.
Let $J\in\mathbb{R}^{m\times n}$ and ${\boldsymbol{w}}=\frac{1}{m}\sum_{i=1}^{m}{\boldsymbol{w}}_{i}$ be the weights defining $\mathcal{A}_{\text{UPGrad}}(J)$ as per (5). Let, as usual, $G=JJ^{\top}$, then,
${\boldsymbol{w}}^{\top}G{\boldsymbol{w}}\varleq{\boldsymbol{1}}^{\top}G{\boldsymbol{w}}.$
Proof.
[…]
$({\boldsymbol{u}}^{\top}G{\boldsymbol{v}})^{2}$ $=\langle{\boldsymbol{u}},{\boldsymbol{v}}\rangle^{2}$
$\varleq\langle{\boldsymbol{u}},{\boldsymbol{u}}\rangle\cdot\langle{\boldsymbol{v}},{\boldsymbol{v}}\rangle$
$={\boldsymbol{u}}^{\top}G{\boldsymbol{u}}\cdot{\boldsymbol{v}}^{\top}G{\boldsymbol{v}}.$
Therefore
${\boldsymbol{w}}^{\top}G{\boldsymbol{w}}$
[…]
$\frac{1}{m^{2}}\sum_{i,j}\sqrt{{\boldsymbol{w}}_{i}^{\top}G{\boldsymbol{w}}_{i}}\cdot\sqrt{{\boldsymbol{w}}_{j}^{\top}G{\boldsymbol{w}}_{j}}$ $\left(\begin{subarray}{c}\text{Cauchy-Schwartz}\\ \text{inequality}\end{subarray}\right)$
[…]
$~{}\sum_{i}\frac{1}{m}\left(\sqrt{{\boldsymbol{w}}_{i}^{\top}G{\boldsymbol{w}}_{i}}\right)^{2}$ (Jensen’s $\left(\begin{subarray}{c}\text{Jensen's inequality}\end{subarray}\right)$
[…]
which concludes the proof. ∎
Lemma 6.
Under the assumptions of Theorem 1, for any ${\boldsymbol{w}}\in{\mathbb{R}}^{m}$,
${\boldsymbol{w}}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{t+1})-{\boldsymbol{f}}({\boldsymbol{x}}_{t}))\varleq\frac{\|{\boldsymbol{w}}\|}{\beta\sqrt{m}}\left(\frac{{\boldsymbol{1}}}{2\sqrt{m}}-\frac{{\boldsymbol{w}}}{\|{\boldsymbol{w}}\|}\right)^{\top}G_{t}{\boldsymbol{w}}_{t}.$
Proof.
For all $t\vargeq 0$, let $J_{t}=\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}_{t})$, $G_{t}=J_{t}J_{t}^{\top}$. Then ${\boldsymbol{x}}_{t+1}={\boldsymbol{x}}_{t}-\eta\mathcal{A}_{\text{UPGrad}}(J_{t})={\boldsymbol{x}}_{t}-\eta J_{t}^{\top}{\boldsymbol{w}}_{t}$. Therefore
${\boldsymbol{w}}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{t+1})-{\boldsymbol{f}}({\boldsymbol{x}}_{t}))$
[…]
which concludes the proof. ∎
Lemma 7.
Under the assumptions of Theorem 1, if ${\boldsymbol{x}}^{\ast}\in X^{\ast}$ is such that ${\boldsymbol{1}}^{\top}{\boldsymbol{f}}({\boldsymbol{x}}^{\ast})\varleq{\boldsymbol{1}}^{\top}{\boldsymbol{f}}({\boldsymbol{x}}_{t})$ for all $t$, then
$\frac{1}{T}\sum_{t=0}^{T-1}{\boldsymbol{w}}_{t}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{t})-{\boldsymbol{f}}({\boldsymbol{x}}^{\ast}))\varleq\frac{1}{T}\bigg{(}{\boldsymbol{1}}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{0})-{\boldsymbol{f}}({\boldsymbol{x}}^{\ast}))+\frac{\beta\sqrt{m}}{2}\|{\boldsymbol{x}}_{0}-{\boldsymbol{x}}^{\ast}\|^{2}\bigg{)}.$ (16)
Proof.
We first bound, for any $t\vargeq 0$, ${\boldsymbol{1}}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{t+1})-{\boldsymbol{f}}({\boldsymbol{x}}_{t}))$ as follows
${\boldsymbol{1}}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{t+1})-{\boldsymbol{f}}({\boldsymbol{x}}_{t}))$
[…]
Summing this over $t=0,\dots,T-1$ yields
$\frac{1}{2\beta\sqrt{m}}\sum_{t=0}^{T-1}{\boldsymbol{w}}_{t}^{\top}G_{t}{\boldsymbol{w}}_{t}$
$\sum_{t=0}^{T-1}{\boldsymbol{1}}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{t})-{\boldsymbol{f}}({\boldsymbol{x}}_{t+1}))$
[…]
${\boldsymbol{1}}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{0})-{\boldsymbol{f}}({\boldsymbol{x}}^{\ast})).$ $\left(\begin{subarray}{c}\text{Assumption}\\ {\boldsymbol{1}}^{\top}{\boldsymbol{f}}({\boldsymbol{x}}^{\ast})\varleq{\boldsymbol{1}}^{\top}{\boldsymbol{f}}({\boldsymbol{x}}_{T})\end{subarray}\right)$ (17)
Since ${\boldsymbol{0}}\preceq{\boldsymbol{w}}_{t}$,
${\boldsymbol{w}}_{t}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{t})-{\boldsymbol{f}}({\boldsymbol{x}}^{\ast}))$
[…]
Summing this over $t=0,\dots,T-1$ yields
$\sum_{t=0}^{T-1}{\boldsymbol{w}}_{t}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{t})-{\boldsymbol{f}}({\boldsymbol{x}}^{\ast}))$
$\frac{1}{2\beta\sqrt{m}}\sum_{t=0}^{\top}{\boldsymbol{w}}_{t}^{\top}G_{t}{\boldsymbol{w}}_{t}+\frac{\beta\sqrt{m}}{2}\bigg{(}\|{\boldsymbol{x}}_{0}-{\boldsymbol{x}}^{\ast}\|^{2}-\|{\boldsymbol{x}}_{T}-{\boldsymbol{x}}^{\ast}\|^{2}\bigg{)}$ (Telescoping $\left(\begin{subarray}{c}\text{Telescoping sum}\end{subarray}\right)$
$\frac{1}{2\beta\sqrt{m}}\sum_{t=0}^{\top}{\boldsymbol{w}}_{t}^{\top}G_{t}{\boldsymbol{w}}_{t}+\frac{\beta\sqrt{m}}{2}\|{\boldsymbol{x}}_{0}-{\boldsymbol{x}}^{\ast}\|^{2}$
[…]
Scaling down this inequality by $T$ yields
$\frac{1}{T}\sum_{t=0}^{T-1}{\boldsymbol{w}}_{t}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{t})-{\boldsymbol{f}}({\boldsymbol{x}}^{\ast}))\varleq\frac{1}{T}\bigg{(}{\boldsymbol{1}}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{0})-{\boldsymbol{f}}({\boldsymbol{x}}^{\ast}))+\frac{\beta\sqrt{m}}{2}\|{\boldsymbol{x}}_{0}-{\boldsymbol{x}}^{\ast}\|^{2}\bigg{)},$
which concludes the proof. ∎
We are now ready to prove Theorem 1.
Proof.
For all $t\vargeq 0$, let $J_{t}=\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}_{t})$, $G_{t}=J_{t}J_{t}^{\top}$. Then
${\boldsymbol{x}}_{t+1}$ $={\boldsymbol{x}}_{t}-\eta\mathcal{A}_{\text{UPGrad}}(J_{t})$
$={\boldsymbol{x}}_{t}-\eta J_{t}^{\top}{\boldsymbol{w}}_{t}.$
Substituting ${\boldsymbol{w}}={\boldsymbol{1}}$ in the term $\frac{{\boldsymbol{1}}}{2\sqrt{m}}-\frac{{\boldsymbol{w}}}{\|{\boldsymbol{w}}\|}$ of Lemma 6 yields
$\frac{{\boldsymbol{1}}}{2\sqrt{m}}-\frac{{\boldsymbol{w}}}{\|{\boldsymbol{w}}\|}$ $=-\frac{{\boldsymbol{1}}}{2\sqrt{m}}$
$\prec{\boldsymbol{0}}.$
Therefore there exists some $\varepsilon>0$ such that any ${\boldsymbol{w}}\in{\mathbb{R}}^{m}$ with $\|{\boldsymbol{1}}-{\boldsymbol{w}}\|<\varepsilon$ satisfies $\frac{{\boldsymbol{1}}}{2\sqrt{m}}\prec\frac{{\boldsymbol{w}}}{\|{\boldsymbol{w}}\|}$. Denote by $B_{\varepsilon}({\boldsymbol{1}})=\{{\boldsymbol{w}}\in{\mathbb{R}}^{m}:\|{\boldsymbol{1}}-{\boldsymbol{w}}\|<\varepsilon\}$, i.e. for all ${\boldsymbol{w}}\in B_{\varepsilon}({\boldsymbol{1}})$, $\frac{{\boldsymbol{1}}}{2\sqrt{m}}\prec\frac{{\boldsymbol{w}}}{\|{\boldsymbol{w}}\|}$. By the non-conflicting property of $\mathcal{A}_{\text{UPGrad}}$, ${\boldsymbol{0}}\preceq G_{t}{\boldsymbol{w}}_{t}$ and therefore for all ${\boldsymbol{w}}\in B_{\varepsilon}({\boldsymbol{1}})$,
${\boldsymbol{w}}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{t+1})-{\boldsymbol{f}}({\boldsymbol{x}}_{t}))$
[…]
$0$
Since ${\boldsymbol{w}}^{\top}{\boldsymbol{f}}({\boldsymbol{x}}_{t})$ is bounded and non-increasing, it converges. Since $B_{\varepsilon}({\boldsymbol{1}})$ contains a basis of ${\mathbb{R}}^{m}$, ${\boldsymbol{f}}({\boldsymbol{x}}_{t})$ converges to some ${\boldsymbol{f}}^{\ast}\in{\mathbb{R}}^{m}$. By assumption on ${\boldsymbol{f}}$, there exists ${\boldsymbol{x}}^{\ast}$ in the Pareto set such that ${\boldsymbol{f}}({\boldsymbol{x}}^{\ast})\preceq{\boldsymbol{f}}^{\ast}$.
We now prove that ${\boldsymbol{f}}({\boldsymbol{x}}^{\ast})={\boldsymbol{f}}^{\ast}$. Since ${\boldsymbol{f}}({\boldsymbol{x}}^{\ast})\preceq{\boldsymbol{f}}^{\ast}$, it is sufficient to show that ${\boldsymbol{1}}^{\top}({\boldsymbol{f}}^{\ast}-{\boldsymbol{f}}({\boldsymbol{x}}^{\ast}))\varleq 0$.
First, the additional assumption of Lemma 7 applies since ${\boldsymbol{1}}^{\top}{\boldsymbol{f}}({\boldsymbol{x}}_{t})$ decreases to ${\boldsymbol{1}}^{\top}{\boldsymbol{f}}^{\ast}$ which is larger than ${\boldsymbol{1}}^{\top}{\boldsymbol{f}}({\boldsymbol{x}}^{\ast})$. Therefore
${\boldsymbol{1}}^{\top}({\boldsymbol{f}}^{\ast}-{\boldsymbol{f}}({\boldsymbol{x}}^{\ast}))$
[…]
Taking the limit as $T\to\infty$, we get
${\boldsymbol{1}}^{\top}({\boldsymbol{f}}^{\ast}-{\boldsymbol{f}}({\boldsymbol{x}}^{\ast}))$
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}{\boldsymbol{w}}_{t}^{\top}({\boldsymbol{f}}^{\ast}-{\boldsymbol{f}}({\boldsymbol{x}}_{t}))$
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\|{\boldsymbol{w}}_{t}\|\cdot\|{\boldsymbol{f}}^{\ast}-{\boldsymbol{f}}({\boldsymbol{x}}_{t})\|$ $\left(\begin{subarray}{c}\text{Cauchy-Schwartz}\\ \text{inequality}\end{subarray}\right)$
[…]
which concludes the proof. ∎
Appendix B Properties of existing aggregators
In the following, we prove the properties of the aggregators from Table 1. Some aggregators, e.g. $\mathcal{A}_{\text{RGW}}$, $\mathcal{A}_{\text{GradDrop}}$ and $\mathcal{A}_{\text{PCGrad}}$, are non-deterministic and are thus not technically functions but rather random variables whose distribution depends on the matrix $J\in\mathbb{R}^{m\times n}$ to aggregate. Still, the properties of Section 2.2 can be easily adapted to a random setting. If $\mathcal{A}$ is a random aggregator, then for any $J$, $\mathcal{A}(J)$ is a random vector in ${\mathbb{R}}^{n}$. The aggregator is non-conflicting if $\mathcal{A}(J)$ is in the dual cone of the rows of $J$ with probability 1. It is linear under scaling if for all $J\in\mathbb{R}^{m\times n}$, there is a – possibly random – matrix $\mathfrak{J}\in\mathbb{R}^{m\times n}$, such that for all ${\boldsymbol{0}}\prec{\boldsymbol{c}}\in{\mathbb{R}}^{m}$, $\mathcal{A}(\operatorname*{diag}({\boldsymbol{c}})\cdot J)=\mathfrak{J}^{\top}\cdot{\boldsymbol{c}}$. Finally, $\mathcal{A}$ is weighted if for any $J\in\mathbb{R}^{m\times n}$ there is, with probability 1, some weighting ${\boldsymbol{w}}\in{\mathbb{R}}^{m}$ such that $\mathcal{A}(J)=J^{\top}\cdot{\boldsymbol{w}}$.
B.1 Mean
$\mathcal{A}_{\text{Mean}}$ simply averages the rows of the input matrix, i.e. for all $J\in\mathbb{R}^{m\times n}$, $\mathcal{A}_{\text{Mean}}(J)=\frac{1}{m}J^{\top}\cdot{\boldsymbol{1}}$. It is weighted with constant weighting equal to $\frac{1}{m}{\boldsymbol{1}}$. For any ${\boldsymbol{c}}\in{\mathbb{R}}^{m}$, $\mathcal{A}_{\text{Mean}}(\operatorname*{diag}({\boldsymbol{c}})\cdot J)=J^{\top}\cdot{\boldsymbol{c}}$, which is linear in ${\boldsymbol{c}}$. $\mathcal{A}_{\text{Mean}}$ is therefore linear under scaling. $\mathcal{A}_{\text{Mean}}\left(\begin{bmatrix}-2\\ 4\end{bmatrix}\right)=\begin{bmatrix}1\end{bmatrix}$, which conflicts with $\begin{bmatrix}-2\end{bmatrix}$, so $\mathcal{A}_{\text{Mean}}$ is not non-conflicting.
B.2 MGDA
The optimization algorithm presented in [6], called MGDA, is tied to a particular method for aggregating the gradients. We thus refer to this aggregator as $\mathcal{A}_{\text{MGDA}}$. The dual problem of this method was also introduced independently in [5]. We show the equivalence between the two solutions to make the analysis of $\mathcal{A}_{\text{MGDA}}$ easier.
Let $J\in\mathbb{R}^{m\times n}$. The aggregation described in [6] is defined by the weighting that is a solution to
$\operatorname*{arg\,min}_{\begin{subarray}{c}{\boldsymbol{0}}\preceq{\boldsymbol{w}}:\\ {\boldsymbol{1}}^{\top}{\boldsymbol{w}}=1\end{subarray}}\|J^{\top}{\boldsymbol{w}}\|^{2}$ (18)
By construction, $\mathcal{A}_{\text{MGDA}}$ is thus weighted.
In Equation (3) of [5], the following problem is studied
$\min_{\begin{subarray}{c}\alpha\in\mathbb{R},{\boldsymbol{x}}\in{\mathbb{R}}^{n}:\\ J{\boldsymbol{x}}\preceq\alpha{\boldsymbol{1}}\end{subarray}}\alpha+\frac{1}{2}\|{\boldsymbol{x}}\|^{2}$ (19)
We show that the problems in (18) and (19) are dual to each other. Furthermore, the duality gap is null since this is a convex problem.
The Lagrangian of the problem in (19) is given by $\mathcal{L}(\alpha,{\boldsymbol{x}},{\boldsymbol{\mu}})=\alpha+\frac{1}{2}\|{\boldsymbol{x}}\|^{2}-{\boldsymbol{\mu}}^{\top}(\alpha{\boldsymbol{1}}-J{\boldsymbol{x}})$. Differentiating w.r.t. $\alpha$ and ${\boldsymbol{x}}$ gives respectively $1-{\boldsymbol{\mu}}^{\top}{\boldsymbol{1}}$ and ${\boldsymbol{x}}+J^{\top}{\boldsymbol{\mu}}$. The dual problem is obtained by setting those two to ${\boldsymbol{0}}$ and then maximizing the Lagrangian on ${\boldsymbol{0}}\preceq{\boldsymbol{\mu}}$ and $\alpha$, i.e.
$\operatorname*{arg\,max}_{\begin{subarray}{c}\alpha,{\boldsymbol{0}}\preceq{\boldsymbol{\mu}}:\\ {\boldsymbol{\mu}}^{\top}{\boldsymbol{1}}=1\end{subarray}}\alpha+\frac{1}{2}\|J^{\top}{\boldsymbol{\mu}}\|^{2}-{\boldsymbol{\mu}}^{\top}(\alpha{\boldsymbol{1}}+JJ^{\top}{\boldsymbol{\mu}})$
[…]
Therefore, (18) and (19) are equivalent, with ${\boldsymbol{x}}=-J^{\top}{\boldsymbol{w}}$.
Observe that since in (19), $\alpha=0$ and ${\boldsymbol{x}}={\boldsymbol{0}}$ is feasible, the objective is non-positive and therefore $\alpha\varleq 0$. Substituting ${\boldsymbol{x}}=-J^{\top}{\boldsymbol{w}}$ in $J\cdot{\boldsymbol{x}}\preceq\alpha{\boldsymbol{1}}\preceq{\boldsymbol{0}}$ yields ${\boldsymbol{0}}\preceq JJ^{\top}{\boldsymbol{w}}$, i.e. ${\boldsymbol{0}}\preceq J\cdot\mathcal{A}_{\text{MGDA}}(J)$, so $\mathcal{A}_{\text{MGDA}}$ is non-conflicting.
With $J=\begin{bmatrix}2&0\\ 0&2\\ a&a\end{bmatrix}$, if $0\varleq a\varleq 1$, $\mathcal{A}_{\text{MGDA}}(J)=\begin{bmatrix}a\\ a\end{bmatrix}$. However, if $a\vargeq 1$, $\mathcal{A}_{\text{MGDA}}(J)=\begin{bmatrix}1\\ 1\end{bmatrix}$. This is not affine in $a$, so $\mathcal{A}_{\text{MGDA}}$ is not linear under scaling. In particular, if any row of $J$ is ${\boldsymbol{0}}$, $\mathcal{A}_{\text{MGDA}}(J)={\boldsymbol{0}}$. This implies that the optimization will stop whenever one objective has converged.
B.3 DualProj
The projection of a gradient of interest onto a dual cone was first described in [33]. When this gradient is the average of the rows of the Jacobian, we call this aggregator $\mathcal{A}_{\text{DualProj}}$. Formally, $\mathcal{A}_{\text{DualProj}}(J)=\frac{1}{m}\cdot\pi_{J}(J^{\top}\cdot{\boldsymbol{1}})$. By construction, $\mathcal{A}_{\text{DualProj}}$ is thus non-conflicting. By Proposition 1, $\mathcal{A}_{\text{DualProj}}(J)=\frac{1}{m}J^{\top}\cdot{\boldsymbol{w}}$, with ${\boldsymbol{w}}\in\operatorname*{arg\,min}_{{\boldsymbol{1}}\preceq{\boldsymbol{v}}}{\boldsymbol{v}}^{\top}JJ^{\top}{\boldsymbol{v}}$. $\mathcal{A}_{\text{DualProj}}$ is thus weighted.
With $J=\begin{bmatrix}2&0\\ -2a&2a\end{bmatrix}$, if $a\vargeq 1$, $\mathcal{A}_{\text{DualProj}}(J)=\begin{bmatrix}0\\ a\end{bmatrix}$. However, if $0.5\varleq a\varleq 1$, $\mathcal{A}_{\text{DualProj}}(J)=\begin{bmatrix}1-a\\ a\end{bmatrix}$. This is not affine in $a$, so $\mathcal{A}_{\text{DualProj}}$ is not linear under scaling.
B.4 PCGrad
$\mathcal{A}_{\text{PCGrad}}$ is described in [11]. It projects each gradient onto the orthogonal hyperplane of other gradients in case of conflict with them, iteratively and in random order. When $m\varleq 2$, $\mathcal{A}_{\text{PCGrad}}$ is deterministic and is such that $\mathcal{A}_{\text{PCGrad}}=m\cdot\mathcal{A}_{\text{UPGrad}}$. Therefore, in this case, it satisfies all three properties. When $m>2$, $\mathcal{A}_{\text{PCGrad}}$ is non-deterministic, so $\mathcal{A}_{\text{PCGrad}}(J)$ is a random vector.
For any index $i\in[m]$, let ${\boldsymbol{g}}_{i}=J^{\top}\cdot{\boldsymbol{e}}_{i}$ and let $P(i)$ be distributed uniformly on the permutations of the elements in $[m]\setminus\{i\}$. For instance, if $m=3$, $P(2)=\begin{bmatrix}1&3\end{bmatrix}^{\top}$ with probability $0.5$ and $P(2)=\begin{bmatrix}3&1\end{bmatrix}^{\top}$ with probability $0.5$. The iterative projection of $\mathcal{A}_{\text{PCGrad}}$ is then defined recursively as:
${\boldsymbol{g}}_{i,1}^{\text{PC}}$ $={\boldsymbol{g}}_{i}$
$j_{k}$ $=P(i)_{k}$
${\boldsymbol{g}}_{i,k+1}^{\text{PC}}$ $={\boldsymbol{g}}_{i,k}^{\text{PC}}-\mathds{1}\{{\boldsymbol{g}}_{i,k}^{\text{PC}}\cdot{\boldsymbol{g}}_{j_{k}}<0\}\frac{{\boldsymbol{g}}_{i,k}^{\text{PC}}\cdot{\boldsymbol{g}}_{j_{k}}}{\|{\boldsymbol{g}}_{j_{k}}\|^{2}}{\boldsymbol{g}}_{j_{k}}$ (20)
We noticed that an equivalent formulation to the conditional projection of (20) is the projection onto the dual cone of $\{{\boldsymbol{g}}_{j_{k}}\}$:
${\boldsymbol{g}}_{i,k+1}^{\text{PC}}=\pi_{{\boldsymbol{g}}_{j_{k}}^{\top}}({\boldsymbol{g}}_{i,k}^{\text{PC}})$ (21)
Finally, the aggregation is given by $\mathcal{A}_{\text{PCGrad}}(J)=\sum_{i=1}^{m}{\boldsymbol{g}}_{i,m}^{\text{PC}}$.
For all $i$, ${\boldsymbol{g}}_{i,m}^{\text{PC}}$ is always a linear combination of rows of $J$. $\mathcal{A}_{\text{PCGrad}}$ is thus weighted.
If $J=\begin{bmatrix}1&0\\ 0&1\\ -0.5&-1\end{bmatrix}$, the only non-conflicting direction is ${\boldsymbol{0}}$. However, $\mathcal{A}_{\text{PCGrad}}(J)$ is uniform over the set $\left\{\begin{bmatrix}0.4\\ 0.2\end{bmatrix},\begin{bmatrix}0.8\\ 0.2\end{bmatrix},\begin{bmatrix}0.4\\ -0.2\end{bmatrix},\begin{bmatrix}0.8\\ -0.2\end{bmatrix}\right\}$. $\mathcal{A}_{\text{PCGrad}}$ is thus not non-conflicting. Note that here, $\mathds{E}[\mathcal{A}_{\text{PCGrad}}(J)]=\begin{bmatrix}0.6&0\end{bmatrix}^{\top}$, so $\mathcal{A}_{\text{PCGrad}}$ is neither non-conflicting in expectation.
To show that $\mathcal{A}_{\text{PCGrad}}$ is linear under scaling, let ${\boldsymbol{0}}\prec{\boldsymbol{c}}\in{\mathbb{R}}^{m}$, ${\boldsymbol{g}}_{i}^{\prime}=c_{i}{\boldsymbol{g}}_{i}$, ${\boldsymbol{g}}_{i,1}^{\prime\text{ PC}}={\boldsymbol{g}}^{\prime}_{i}$ and ${\boldsymbol{g}}_{i,k+1}^{\prime\text{ PC}}=\pi_{{\boldsymbol{g}}_{j_{k}}^{\prime\top}}({\boldsymbol{g}}_{i,k}^{\prime\text{ PC}})$. We show by induction that ${\boldsymbol{g}}_{i,k}^{\prime\text{ PC}}=c_{i}{\boldsymbol{g}}_{i,k}^{\text{PC}}$.
The base case is given by ${\boldsymbol{g}}_{i,1}^{\prime\text{ PC}}={\boldsymbol{g}}_{i}^{\prime}=c_{i}{\boldsymbol{g}}_{i}=c_{i}{\boldsymbol{g}}_{i,1}^{\text{PC}}$.
Then, assuming the induction hypothesis ${\boldsymbol{g}}_{i,k}^{\prime\text{ PC}}=c_{i}{\boldsymbol{g}}_{i,k}^{\text{PC}}$, we show ${\boldsymbol{g}}_{i,k+1}^{\prime\text{ PC}}=c_{i}{\boldsymbol{g}}_{i,k+1}^{\text{PC}}$:
${\boldsymbol{g}}_{i,k+1}^{\prime\text{ PC}}=\pi_{c_{j_{k}}{\boldsymbol{g}}_{j_{k}}^{\top}}(c_{i}{\boldsymbol{g}}_{i,k}^{\text{ PC}})$ (Induction $\left(\begin{subarray}{c}\text{Induction hypothesis}\end{subarray}\right)$
${\boldsymbol{g}}_{i,k+1}^{\prime\text{ PC}}=c_{i}\pi_{{\boldsymbol{g}}_{j_{k}}^{\top}}({\boldsymbol{g}}_{i,k}^{\text{ PC}})$ $\left(\begin{subarray}{c}0<c_{i}\text{ and }0<c_{j_{k}}\end{subarray}\right)$
[…]
Therefore $\mathcal{A}_{\text{PCGrad}}(\operatorname*{diag}({\boldsymbol{c}})\cdot J)=\sum_{i=1}^{m}c_{i}{\boldsymbol{g}}_{i,m}^{\text{PC}}$, so $\mathcal{A}_{\text{PCGrad}}$ is linear under scaling.
B.5 GradDrop
The aggregator used by the GradDrop layer, which we denote $\mathcal{A}_{\text{GradDrop}}$, is described in [17]. It is non-deterministic, so $\mathcal{A}_{\text{GradDrop}}(J)$ is a random vector. Given $J\in\mathbb{R}^{m\times n}$, let $|J|\in\mathbb{R}^{m\times n}$ be the element-wise absolute value of $J$. Let $P\in{\mathbb{R}}^{n}$ be such that $P=\frac{1}{2}\left({\boldsymbol{1}}+\frac{J^{\top}\cdot{\boldsymbol{1}}}{|J|^{\top}\cdot{\boldsymbol{1}}}\right)$, where the division is element-wise. Each coordinate $i\in[n]$ is independently assigned to the set $A$ with probability $P_{i}$ and to the set $B$ otherwise. The aggregation at coordinate $i\in A$ is given by the sum of all positive $J_{ji}$, for $j\in[m]$. The aggregation at coordinate $i\in B$ is given by the sum of all negative $J_{ji}$, for $j\in[m]$.
If $J=\begin{bmatrix}-2\\ 1\end{bmatrix}$, then $P=\begin{bmatrix}1/3\end{bmatrix}$. Therefore, $\mathds{P}[\mathcal{A}_{\text{GradDrop}}(J)=\begin{bmatrix}-2\end{bmatrix}]=2/3$ and $\mathds{P}[\mathcal{A}_{\text{GradDrop}}(J)=\begin{bmatrix}1\end{bmatrix}]=1/3$. Therefore, $\mathcal{A}_{\text{GradDrop}}$ is not non-conflicting (even in expectation).
$\mathcal{A}_{\text{GradDrop}}$ is not linear under scaling. Indeed, suppose that $J=\begin{bmatrix}1&-1\\ -1&1\end{bmatrix}$ then $P=\frac{1}{2}\cdot{\boldsymbol{1}}$ and the aggregation is one of the four vectors [±$\begin{bmatrix}\pm 1&\pm 1\end{bmatrix}^{\top}$ with equal probability. Scaling the first line of $J$ by $2$ yields $J=\begin{bmatrix}2&-2\\ -1&1\end{bmatrix}$, $P=\begin{bmatrix}2/3&1/3\end{bmatrix}^{\top}$ which cannot lead to a uniform distribution over four elements.
With $J=\begin{bmatrix}1&-1\\ -1&1\end{bmatrix}$, the span of $J$ does not include $\begin{bmatrix}1&1\end{bmatrix}^{\top}$ nor $\begin{bmatrix}-1&-1\end{bmatrix}^{\top}$. Therefore, $\mathcal{A}_{\text{GradDrop}}$ is not weighted.
B.6 IMTL-G
In [13], the authors describe a method to impartially balance gradients by weighting them. Let ${\boldsymbol{g}}_{i}$ be the $i$’th row of $J$ and let ${\boldsymbol{u}}_{i}=\frac{{\boldsymbol{g}}_{i}}{\|{\boldsymbol{g}}_{i}\|}$. Then, they want to find a combination ${\boldsymbol{g}}=\sum_{i=1}^{m}\alpha_{i}{\boldsymbol{g}}_{i}$ such that ${\boldsymbol{g}}^{\top}{\boldsymbol{u}}_{i}$ is equal for all $i$. Let $U=\begin{bmatrix}{\boldsymbol{u}}_{1}-{\boldsymbol{u}}_{2}&\dots&{\boldsymbol{u}}_{1}-{\boldsymbol{u}}_{m}\end{bmatrix}^{\top}$, $D=\begin{bmatrix}{\boldsymbol{g}}_{1}-{\boldsymbol{g}}_{2}&\dots&{\boldsymbol{g}}_{1}-{\boldsymbol{g}}_{m}\end{bmatrix}^{\top}$. If ${\boldsymbol{\alpha}}_{2:m}=\begin{bmatrix}\alpha_{2}&\dots&\alpha_{m}\end{bmatrix}^{\top}$, then ${\boldsymbol{\alpha}}_{2:m}=(UD^{\top})^{-1}U\cdot{\boldsymbol{g}}_{1}$ and $\alpha_{1}=1-\sum_{i=2}^{m}\alpha_{i}$. Notice that this is defined only when the gradients are linearly independent. Strictly speaking, this is thus not an aggregator since it can only be computed on matrices of rank $m$. To generalize to the case where $J$ has rank less than $m$, we provide what we believe to be a more natural formulation, which is equivalent in the case where $J$ has rank $m$. Let $J\in\mathbb{R}^{m\times n}$ and $J^{\prime}\in\mathbb{R}^{m\times n}$ be the row normalized version of $J$. Reformulating the desired property that ${\boldsymbol{g}}^{\top}{\boldsymbol{u}}_{i}$ is constant over $i$, in our framework, we solve for the following objective. Let ${\boldsymbol{w}}\in{\mathbb{R}}^{m}$ be such that ${\boldsymbol{1}}^{\top}{\boldsymbol{w}}=1$ and $J^{\prime}J^{\top}{\boldsymbol{w}}~{}\propto~{}{\boldsymbol{1}}$. The aggregated vector is then $J^{\top}{\boldsymbol{w}}$. If $J^{\prime}J^{\top}$ is full rank, then ${\boldsymbol{w}}~{}\propto~{}(J^{\prime}J^{\top})^{-1}{\boldsymbol{1}}$. Letting ${\boldsymbol{d}}\in{\mathbb{R}}^{m}$ be the vector of norms of the rows of $J$, this can be rewritten as ${\boldsymbol{w}}~{}\propto~{}(JJ^{\top})^{-1}{\boldsymbol{d}}$. When $JJ^{\top}$ is not full rank, using the Moore-Penrose inverse instead of the usual inverse seems to be a reasonable generalization. The vector of weights is then given by ${\boldsymbol{w}}~{}\propto~{}(JJ^{\top})^{\dagger}{\boldsymbol{d}}$. Finally whenever ${\boldsymbol{1}}^{\top}(JJ^{\top})^{\dagger}{\boldsymbol{d}}=0$, set ${\boldsymbol{w}}={\boldsymbol{0}}$. Throughout this paper, $\mathcal{A}_{\text{IMTL-G}}$ refers to this generalization.
Clearly, $\mathcal{A}_{\text{IMTL-G}}$ is weighted. We now show that $\mathcal{A}_{\text{IMTL-G}}$ is not non-conflicting. If $J=\begin{bmatrix}1&-1&-1\end{bmatrix}^{\top}$, then ${\boldsymbol{d}}=\begin{bmatrix}1&1&1\end{bmatrix}^{\top}$. Also $JJ^{\top}=\begin{bmatrix}1&-1&-1\\ -1&1&1\\ -1&1&1\end{bmatrix}$ and therefore $(JJ^{\top})^{\dagger}\cdot{\boldsymbol{d}}=\frac{1}{9}\begin{bmatrix}-1&1&1\end{bmatrix}^{\top}$. Note that in order to ensure ${\boldsymbol{1}}^{\top}{\boldsymbol{w}}=1$, we have to scale by ${\boldsymbol{1}}^{\top}(JJ^{\top})^{\dagger}\cdot{\boldsymbol{d}}=\frac{1}{9}$. Therefore ${\boldsymbol{w}}=\begin{bmatrix}-1&1&1\end{bmatrix}^{\top}$, $\mathcal{A}_{\text{IMTL-G}}(J)=J^{\top}{\boldsymbol{w}}=\begin{bmatrix}-3\end{bmatrix}$ and $J\cdot\mathcal{A}_{\text{IMTL-G}}(J)=\begin{bmatrix}-3&3&3\end{bmatrix}^{\top}$. $\mathcal{A}_{\text{IMTL-G}}$ is thus not non-conflicting.
It should be noted that when $J$ has rank $m$, $\mathcal{A}_{\text{IMTL-G}}$ seems to be non-conflicting. Thus, it would be possible to make a different non-conflicting generalization, for instance, by deciding ${\boldsymbol{0}}$ when the matrix is not full rank.
$\mathcal{A}_{\text{IMTL-G}}$ is not linear under scaling. if ${\boldsymbol{0}}\prec{\boldsymbol{c}}\in{\mathbb{R}}^{m}$, then the weights associated to $\operatorname*{diag}({\boldsymbol{c}})\cdot J$ are proportional to $(\operatorname*{diag}({\boldsymbol{c}})\cdot JJ^{\top}\cdot\operatorname*{diag}({\boldsymbol{c}}))^{\dagger}\cdot\operatorname*{diag}({\boldsymbol{c}})\cdot{\boldsymbol{d}}=\operatorname*{diag}({\boldsymbol{c}})^{\dagger}\cdot(JJ^{\top})^{\dagger}\cdot{\boldsymbol{d}}$. The normalization value is therefore $\frac{1}{{\boldsymbol{1}}^{\top}\operatorname*{diag}({\boldsymbol{c}})^{\dagger}\cdot(JJ^{\top})^{\dagger}\cdot{\boldsymbol{d}}}$. Therefore
$\mathcal{A}_{\text{IMTL-G}}(\operatorname*{diag}({\boldsymbol{c}})\cdot J)$ $=\frac{J^{\top}\cdot\operatorname*{diag}({\boldsymbol{c}})\cdot\operatorname*{diag}({\boldsymbol{c}})^{\dagger}\cdot(JJ^{\top})^{\dagger}\cdot{\boldsymbol{d}}}{{\boldsymbol{1}}^{\top}\operatorname*{diag}({\boldsymbol{c}})^{\dagger}\cdot(JJ^{\top})^{\dagger}\cdot{\boldsymbol{d}}}$
$=\frac{J^{\top}(JJ^{\top})^{\dagger}\cdot{\boldsymbol{d}}}{{\boldsymbol{1}}^{\top}\operatorname*{diag}({\boldsymbol{c}})^{\dagger}\cdot(JJ^{\top})^{\dagger}\cdot{\boldsymbol{d}}}$
This is not linear in ${\boldsymbol{c}}$, so $\mathcal{A}_{\text{IMTL-G}}$ is not linear under scaling.
B.7 CAGrad
$\mathcal{A}_{\text{CAGrad}}$ is described in [12]. It is parameterized by $c\in\left[0,1\right[{}$. If $c=0$, this is equivalent to $\mathcal{A}_{\text{Mean}}$. Therefore, we restrict our analysis to the case $c>0$. For any $J\in\mathbb{R}^{m\times n}$, let $\bar{\boldsymbol{g}}$ be the average gradient $\frac{1}{m}J^{\top}\cdot{\boldsymbol{1}}$. The aggregated vector is defined as the minimizer $d\in{\mathbb{R}}^{n}$ of the following optimization problem
$\operatorname*{arg\,max}_{\begin{subarray}{c}{\boldsymbol{d}}\in{\mathbb{R}}^{n}:\\ \|{\boldsymbol{d}}-\overline{\boldsymbol{g}}\|\varleq c\|\overline{\boldsymbol{g}}\|\end{subarray}}\min_{i\in[m]}{\boldsymbol{e}}_{i}^{\top}J{\boldsymbol{d}}$
For $10<c<$ 1, $\mathcal{A}_{\text{CAGrad}}$ can have conflict. If $J=\begin{bmatrix}2&0\\ -2a-2&2\end{bmatrix}$, then $\overline{\boldsymbol{g}}=\begin{bmatrix}-a&1\end{bmatrix}^{\top}$ and $\|\overline{\boldsymbol{g}}\|=\sqrt{a^{2}+1}$. Observe that any ${\boldsymbol{d}}\in{\mathbb{R}}^{n}$ satisfying the constraint $\|{\boldsymbol{d}}-\overline{\boldsymbol{g}}\|\varleq c\|\overline{\boldsymbol{g}}\|$ has first coordinate at most $-a+c\sqrt{a^{2}+1}$. Suppose that $a$ is such that $-a+c\sqrt{a^{2}+1}<0$, i.e. $\sqrt{\frac{c^{2}}{1-c^{2}}}<a$. For such an $a$, any feasible $d$ has a negative first coordinate, making it conflict with the first row of $J$. $\mathcal{A}_{\text{CAGrad}}$ is thus not non-conflicting.
Note that if we generalize to the case $c\vargeq 1$, as suggested in the original paper, then ${\boldsymbol{d}}={\boldsymbol{0}}$ becomes feasible, which yields $\min_{i\in[m]}{\boldsymbol{e}}_{i}^{\top}J{\boldsymbol{d}}=0$. Therefore the optimal ${\boldsymbol{d}}$ is such that $0\varleq\min_{i\in[m]}{\boldsymbol{e}}_{i}^{\top}J{\boldsymbol{d}}$, i.e. ${\boldsymbol{0}}\preceq J{\boldsymbol{d}}$. With $c\vargeq 1$, $\mathcal{A}_{\text{CAGrad}}$ would thus be non-conflicting.
In [12], they formulate $\mathcal{A}_{\text{CAGrad}}$ using its dual formulation. If ${\boldsymbol{w}}\in{\mathbb{R}}^{m}$ is a solution to
$\min_{\begin{subarray}{c}{\boldsymbol{0}}\preceq{\boldsymbol{w}}:\\ {\boldsymbol{1}}^{\top}{\boldsymbol{w}}=1\end{subarray}}{\boldsymbol{1}}^{\top}JJ^{\top}{\boldsymbol{w}}+c\cdot\|J^{\top}{\boldsymbol{1}}\|\cdot\|J^{\top}{\boldsymbol{w}}\|$
then the aggregation is $\frac{1}{m}J^{\top}\left({\boldsymbol{1}}+\frac{c\|J^{\top}{\boldsymbol{1}}\|}{\|J^{\top}{\boldsymbol{w}}\|}{\boldsymbol{w}}\right)$. Therefore, $\mathcal{A}_{\text{CAGrad}}$ is weighted.
$\mathcal{A}_{\text{CAGrad}}$ is not linear under scaling. We sketch a proof of this. Suppose that $J=\begin{bmatrix}2&0\\ 0&2a\end{bmatrix}$. Note that $\overline{\boldsymbol{g}}=\begin{bmatrix}1&a\end{bmatrix}^{\top}$ and $\|\overline{\boldsymbol{g}}\|=\sqrt{1+a^{2}}$. One can show that the constraint $\|{\boldsymbol{d}}-\overline{\boldsymbol{g}}\|\varleq c\|\overline{\boldsymbol{g}}\|$ needs to be satisfied with equality since, otherwise, we can scale ${\boldsymbol{d}}$ to make the objective larger. Substituting $J$ in $\min_{i\in[m]}{\boldsymbol{e}}_{i}^{\top}J{\boldsymbol{d}}$ yields $2\min(d_{1},ad_{2})$. For any $a$ such that $c\sqrt{1+a^{2}}+1<a^{2}$, it can be shown that the optimal ${\boldsymbol{d}}$ is such that $d_{1}<ad_{2}$. In that case the inner minimum over $i$ is $2d_{1}$ and, to satisfy $\|{\boldsymbol{d}}-\overline{\boldsymbol{g}}\|=c\|\overline{\boldsymbol{g}}\|$, the KKT conditions over the Lagrangian yield ${\boldsymbol{d}}-\overline{\boldsymbol{g}}~{}\propto~{}\nabla_{\boldsymbol{d}}d_{1}=\begin{bmatrix}1&0\end{bmatrix}^{\top}$. This yields ${\boldsymbol{d}}=\begin{bmatrix}c\cdot\|\overline{\boldsymbol{g}}\|+1\\ a\end{bmatrix}=\begin{bmatrix}c\sqrt{1+a^{2}}+1\\ a\end{bmatrix}$. This is not affine in $a$; therefore, $\mathcal{A}_{\text{CAGrad}}$ is not linear under scaling.
B.8 RGW
$\mathcal{A}_{\text{RGW}}$ is defined in [14] as the weighted sum of the rows of the input matrix, with a random weighting. The weighting is obtained by sampling $m$ i.i.d. normally distributed random variables and applying a softmax. Formally, $\mathcal{A}_{\text{RGW}}(J)=J^{\top}\cdot\sigma(W)$, with $W\sim N({\boldsymbol{0}},I)$. By design, $\mathcal{A}_{\text{RGW}}$ is thus weighted.
When $J=\begin{bmatrix}1\\ -1\end{bmatrix}$, the only non-conflicting solution is ${\boldsymbol{0}}$. However, $\mathds{P}[\mathcal{A}_{\text{RGW}}(J)={\boldsymbol{0}}]=0$. $\mathcal{A}_{\text{RGW}}$ is thus not non-conflicting.
$\mathcal{A}_{\text{RGW}}(\operatorname*{diag}({\boldsymbol{c}})\cdot J)=(\operatorname*{diag}({\boldsymbol{c}})\cdot J)^{\top}\cdot\sigma(W)=J^{\top}\cdot\operatorname*{diag}({\boldsymbol{c}})\cdot\sigma(W)=J^{\top}\cdot\operatorname*{diag}(\sigma(W))\cdot{\boldsymbol{c}}$. Therefore, $\mathcal{A}_{\text{RGW}}$ is linear under scaling in distribution.
B.9 Nash-MTL
Nash-MTL is described in [15]. Unfortunately, we were not able to verify the proof of Claim 3.1, and we believe that the official implementation of Nash-MTL may mismatch the desired objective by which it is defined. Therefore, we only analyze the initial objective even though our experiments for this aggregator are conducted with the official implementation.
Let $J\in\mathbb{R}^{m\times n}$ and $\varepsilon>0$. Let also $B_{\varepsilon}=\{{\boldsymbol{d}}\in{\mathbb{R}}^{n}:\|{\boldsymbol{d}}\|\varleq\varepsilon,{\boldsymbol{0}}\preceq J{\boldsymbol{d}}\}$. $\mathcal{A}_{\text{Nash-MTL}}$ is then defined as
$\mathcal{A}_{\text{Nash-MTL}}(J)$ $\in\operatorname*{arg\,max}_{{\boldsymbol{d}}\in B_{\varepsilon}}\sum_{i=1}^{m}\log({\boldsymbol{e}}_{i}^{\top}J{\boldsymbol{d}})$
By the constraint, $\mathcal{A}_{\text{Nash-MTL}}$ is non-conflicting. If an aggregator $\mathcal{A}$ is linear under scaling, it should be the case that $\mathcal{A}(aJ)=a\mathcal{A}(J)$ for any scalar $a>0$ and any $J\in\mathbb{R}^{m\times n}$. However, $\log(a{\boldsymbol{e}}_{i}^{\top}J{\boldsymbol{d}})=\log({\boldsymbol{e}}_{i}^{\top}J{\boldsymbol{d}})+\log(a)$. This means that scaling by a scalar does not impact aggregation. Since this is not the trivial ${\boldsymbol{0}}$ aggregator, $\mathcal{A}_{\text{Nash-MTL}}$ is not linear under scaling.
If ${\boldsymbol{d}}^{\prime}$ is the projection of ${\boldsymbol{d}}$ onto the span of the rows of $J$, then $\|{\boldsymbol{d}}^{\prime}\|\varleq\|{\boldsymbol{d}}\|$ and $J{\boldsymbol{d}}=J{\boldsymbol{d}}^{\prime}$. Therefore, ${\boldsymbol{d}}^{\prime}$ is feasible and has the same associated objective value. Therefore, without loss of generality, $\mathcal{A}_{\text{Nash-MTL}}$ is weighted.
B.10 Aligned-MTL
The Aligned-MTL method for balancing the Jacobian is described in [16]. For simplicity, we fix the vector of preferences to $\frac{1}{m}{\boldsymbol{1}}$, but the proofs can be adapted for any non-trivial vector. Given $J\in\mathbb{R}^{m\times n}$, compute the eigen-decomposition of $G=JJ^{\top}=V\Sigma^{2}V^{\top}$. Let $\Sigma^{\dagger}$ be the diagonal matrix whose non-zero elements are the inverse of corresponding non-zero diagonal elements of $\Sigma$. Let $\sigma_{\min}=\min_{i\in[m],\Sigma_{ii}\neq 0}\Sigma_{ii}$. Then the aggregation by $\mathcal{A}_{\text{Aligned-MTL}}$ is equal to $\frac{1}{m}J^{\top}\cdot{\boldsymbol{w}}$, with ${\boldsymbol{w}}=\sigma_{\min}\cdot V\Sigma^{\dagger}V^{\top}\cdot{\boldsymbol{1}}$. It follows immediately that $\mathcal{A}_{\text{Aligned-MTL}}$ is weighted.
If the SVD of $J$ is $V\Sigma U^{\top}$, then $J^{\top}{\boldsymbol{w}}=\sigma_{\min}UPV^{\top}{\boldsymbol{1}}$ with $P=\Sigma^{\dagger}\Sigma$ a diagonal projection matrix with $1$s corresponding to non zero elements of $\Sigma$ and 00s everywhere else. Further, $J\cdot\mathcal{A}_{\text{Aligned-MTL}}(J)=\frac{\sigma_{\min}}{m}\cdot V\Sigma V^{\top}{\boldsymbol{1}}$. Taking $V=\frac{1}{2}\begin{bmatrix}\sqrt{3}&1\\ -1&\sqrt{3}\end{bmatrix}$, and $\Sigma=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}$ yields $J\cdot\mathcal{A}_{\text{Aligned-MTL}}(J)=\frac{1}{2}\cdot V\Sigma V^{\top}{\boldsymbol{1}}=\frac{1}{8}\begin{bmatrix}-3+\sqrt{3}&1-\sqrt{3}\end{bmatrix}^{\top}$ which is not non-negative. $\mathcal{A}_{\text{Aligned-MTL}}$ is thus not non-conflicting.
If $J=\begin{bmatrix}1&0\\ 0&a\end{bmatrix}$, then $U=V=I$, $\Sigma=J$. For $0<a\varleq 1$, $\sigma_{\min}=a$, therefore $\mathcal{A}_{\text{Aligned-MTL}}(J)=\frac{a}{2}\cdot{\boldsymbol{1}}$. For $1<a$, $\sigma_{\min}=1$ and therefore $\mathcal{A}_{\text{Aligned-MTL}}(J)=\frac{{\boldsymbol{1}}}{2}$, which makes $\mathcal{A}_{\text{Aligned-MTL}}$ not linear under scaling.
Appendix C Experimental settings
For all of our experiments, we used PyTorch [55]. We have developed an open-source library [^4] Available at https://github.com/TorchJD/torchjd on top of it to enable Jacobian descent easily. This library is designed to be reusable for many other use cases than the experiments presented in our work. To separate them from the library, the experiments have been conducted with a different code repository [^5] Available at https://github.com/ValerianRey/jde mainly using PyTorch and our library.
C.1 Learning rate selection
The learning rate has a very important impact on the speed of optimization. To make the comparisons as fair as possible, we always show the results corresponding to the best learning rate. We have selected the area under the loss curve as the criterion to compare the learning rates. This choice is arbitrary but seems to work well in practice: lower area under the loss curve means that the optimization is fast (quick loss decrease) and stable (few bumps in the loss curve). Concretely, for each random rerun and for each aggregator, we first try 22 learning rates from $10^{-5}$ to $10^{2}$, increasing by a factor $10^{\frac{1}{3}}$ every time. The two best learning rates from this range then define a refined range of plausible good learning rates, going from the smallest of those two multiplied by $10^{-\frac{1}{3}}$ to the largest of those two multiplied by $10^{\frac{1}{3}}$. This margin makes it unlikely for the best learning rate to lie out of the refined range. After this, 50 learning rates from the refined range are tried. These learning rates are evenly spaced in the exponent domain. The one with the best area under the loss curve is then selected and presented in the plots. Note that for simplicity, we have always used a constant learning rate, i.e. no learning rate scheduler was used.
We think that this approach has the advantage of being quite simple and very precise, thus giving trustworthy results. However, it requires a total of 72 trainings for each aggregator, each random rerun and each dataset. For this reason, we have opted for working on small subsets of the original datasets.
C.2 Random reruns and standard error of the mean
To get an idea of confidence in our results, every experiment is performed 8 times on a different seed and on a different subset, of size 1024, of the training dataset. The seed used for run $i\in[8]$ is always simply set to $i$. Because each random rerun includes the full learning rate selection method described in Appendix C.1, it is sensible to consider the 8 sets of results as i.i.d. For each point of both the loss curves and the cosine similarity curves, we thus compute the estimated standard error of the mean with the usual formula $\frac{1}{\sqrt{8}}\sqrt{\frac{\sum_{i=1}^{8}(v_{i}-\bar{v})^{2}}{8-1}}$, where $v_{i}$ is the value of a point of the curve for random rerun $i$, and $\bar{v}$ is the average value of this point over the 8 runs.
C.3 Model architectures
In all experiments, the models are simple convolutional neural networks. All convolutions always have a stride of 1x1, a kernel size of 3x3, a learnable bias and no padding. All linear layers always have a learnable bias. The activation function is the exponential linear unit [56]. The full architectures are given in Tables 234 and 5. Note that these architectures have been fixed arbitrarily, i.e. they were not optimized through some hyper-parameter selection. The weights of the model have been initialized with the default initialization scheme of PyTorch.
Table 2: Architecture used for SVHN
Conv2d (3 input channels, 16 output channels, 1 groups), ELU
Conv2d (16 input channels, 32 output channels, 16 groups)
MaxPool2d (stride of 2x2, kernel size of 2x2), ELU
Conv2d (32 input channels, 32 output channels, 32 groups)
MaxPool2d (stride of 3x3, kernel size of 3x3), ELU, Flatten
Linear (512 input features, 64 output features), ELU
Linear (64 input features, 10 outputs)
Table 3: Architecture used for CIFAR-10
Conv2d (3 input channels, 32 output channels, 1 group), ELU
Conv2d (32 input channels, 64 output channels, 32 groups)
MaxPool2d (stride of 2x2, kernel size of 2x2), ELU
Conv2d (64 input channels, 64 output channels, 64 groups)
MaxPool2d (stride of 3x3, kernel size of 3x3), ELU, Flatten
Linear (1024 input features, 128 output features), ELU
Linear (128 input features, 10 outputs)
Table 4: Architecture used for EuroSAT
Conv2d (3 input channels, 32 output channels, 1 group)
MaxPool2d (stride of 2x2, kernel size of 2x2), ELU
Conv2d (32 input channels, 64 output channels, 32 groups)
MaxPool2d (stride of 2x2, kernel size of 2x2), ELU
Conv2d (64 input channels, 64 output channels, 64 groups)
MaxPool2d (stride of 3x3, kernel size of 3x3), ELU, Flatten
Linear (1024 input features, 128 output features), ELU
Linear (128 input features, 10 outputs)
Table 5: Architecture used for MNIST, Fashion-MNIST and Kuzushiji-MNIST
Conv2d (1 input channel, 32 output channels, 1 group), ELU
Conv2d (32 input channels, 64 output channels, 1 group)
MaxPool2d (stride of 2x2, kernel size of 2x2), ELU
Conv2d (64 input channels, 64 output channels, 1 group)
MaxPool2d (stride of 3x3, kernel size of 3x3), ELU, Flatten
Linear (576 input features, 128 output features), ELU
Linear (128 input features, 10 outputs)
C.4 Optimizer
For all experiments except those described in Appendix D.3, we always use the basic SGD optimizer of PyTorch, without any regularization or momentum. Note that here, SGD refers to the PyTorch optimizer that updates the parameters of the model in the opposite direction of the gradient, which, in our case, is replaced by the aggregation of the Jacobian matrix. In the rest of this paper, SGD refers to the whole stochastic gradient descent algorithm. In the experiments of Appendix D.3, we use Adam instead, to study its interactions with JD.
C.5 Loss function
The loss function is always the usual cross-entropy, with the default parameters of PyTorch.
C.6 Preprocessing
The inputs are always normalized per-channel based on the mean and standard deviation computed on the entire training split of the dataset.
C.7 Iterations and computational budget
The numbers of epochs and the corresponding numbers of iterations for all datasets are provided in Table 6, along with the required number of NVIDIA L4 GPU-hours, to run all 72 learning rates for the 11 aggregators, on a single seed. The total computational budget to run the main experiments on 8 seeds was thus around 760 GPU-hours. Additionally, we have used a total of about 100 GPU-hours for the experiments varying the batch size and using Adam, and about 200 more GPU-hours were used for early investigations.
Table 6: Numbers of epochs, iterations and GPU-hours for each dataset
Dataset Epochs Iterations GPU-Hours
SVHN 25 800 17
CIFAR-10 20 640 15
EuroSAT 30 960 32
MNIST 8 256 6
Fashion-MNIST 25 800 17
Kuzushiji-MNIST 10 320 8
Appendix D Additional experimental results
In this appendix, we provide additional experimental results about IWRM.
D.1 All datasets and all aggregators
Figures 3, 4, 5, 6, 7 and 8 show the full results of the experiments described in Section 5 on SVHN, CIFAR-10, EuroSAT, MNIST, Fashion-MNIST and Kuzushiji-MNIST, respectively. For readability, the results are displayed on three different plots for each dataset. We always show $\mathcal{A}_{\text{UPGrad}}$ and $\mathcal{A}_{\text{Mean}}$ for reference. The exact experimental settings are described in Appendix C.
It should be noted that some of these aggregators were not developed as general-purpose aggregators, but mostly for the use case of multi-task learning, with one gradient per task. Our experiments present a more challenging setting than multi-task learning optimization, because conflict between rows of the Jacobian is typically higher. Besides, for some aggregators, e.g. $\mathcal{A}_{\text{GradDrop}}$ and $\mathcal{A}_{\text{IMTL-G}}$, it was advised to make the aggregation of gradients w.r.t. an internal activation (such as the last shared representation), rather than w.r.t. the parameters of the model [17, 13]. To enable comparison, we instead always aggregated the Jacobian w.r.t. all parameters.
We can see that $\mathcal{A}_{\text{UPGrad}}$ provides a significant improvement over $\mathcal{A}_{\text{Mean}}$ on all datasets. Moreover, the performance gaps seem to be linked to the difficulty of the dataset, which suggests that experimenting with harder tasks is a promising future direction. The intrinsic randomness of $\mathcal{A}_{\text{RGW}}$ and $\mathcal{A}_{\text{GradDrop}}$ reduces the train set performance, but it could positively impact the generalization, which we do not study here. We suspect the disappointing results of $\mathcal{A}_{\text{Nash-MTL}}$ to be caused by issues in the official implementation that we used, leading to instability.
(a) Training loss
(b) Update similarity to the SGD update
(c) Training loss
(d) Update similarity to the SGD update
(e) Training loss
(f) Update similarity to the SGD update
Figure 3: SVHN results.
(a) Training loss
(b) Update similarity to the SGD update
(c) Training loss
(d) Update similarity to the SGD update
(e) Training loss
(f) Update similarity to the SGD update
Figure 4: CIFAR-10 results.
(a) Training loss
(b) Update similarity to the SGD update
(c) Training loss
(d) Update similarity to the SGD update
(e) Training loss
(f) Update similarity to the SGD update
Figure 5: EuroSAT results.
(a) Training loss
(b) Update similarity to the SGD update
(c) Training loss
(d) Update similarity to the SGD update
(e) Training loss
(f) Update similarity to the SGD update
Figure 6: MNIST results.
(a) Training loss
(b) Update similarity to the SGD update
(c) Training loss
(d) Update similarity to the SGD update
(e) Training loss
(f) Update similarity to the SGD update
Figure 7: Fashion-MNIST results.
(a) Training loss
(b) Update similarity to the SGD update
(c) Training loss
(d) Update similarity to the SGD update
(e) Training loss
(f) Update similarity to the SGD update
Figure 8: Kuzushiji-MNIST results.
D.2 Varying the batch size
Figure 9 shows the results on CIFAR-10 with $\mathcal{A}_{\text{UPGrad}}$ when varying the batch size from 4 to 64. Concretely, because we are using SSJD, this makes the number of rows of the sub-Jacobian aggregated at each step vary from 4 to 64. Recall that IWRM with SSJD and $\mathcal{A}_{\text{Mean}}$ is equivalent to ERM with SGD. We see that with a small batch size, in this setting, $\mathcal{A}_{\text{UPGrad}}$ becomes very similar to $\mathcal{A}_{\text{Mean}}$. This is not surprising, since with a batch size of 1, both would be equivalent. Conversely, a larger batch size increases the gap between $\mathcal{A}_{\text{UPGrad}}$ and $\mathcal{A}_{\text{Mean}}$. Since the projections of $\mathcal{A}_{\text{UPGrad}}$ are onto the dual cone of more rows, each step becomes non-conflicting with respect to more of the original 1024 objectives, pushing even further the benefits of the non-conflicting property. In other words, increasing the batch size refines the dual cone, thereby improving the quality of the projections. It would be interesting to study more theoretically the impact of batch size in this setting.
(a) $\text{BS}=4$: Training loss
(b) $\text{BS}=4$: Update similarity to the SGD update
(c) $\text{BS}=16$: Training loss
(d) $\text{BS}=16$: Update similarity to the SGD update
(e) $\text{BS}=64$: Training loss
(f) $\text{BS}=64$: Update similarity to the SGD update
Figure 9: CIFAR-10 results with different batch sizes (BS). The number of epochs is always 20, so the number of iterations varies.
D.3 Compatibility with Adam
Figure 10 gives the results on CIFAR-10 and SVHN when using Adam rather than the SGD optimizer. Concretely, this corresponds to the Adam algorithm in which the gradient is replaced by the aggregation of the Jacobian. The learning rate is still tuned as described in Appendix C.1, but the other hyperparameters of Adam are fixed to the default values of PyTorch, i.e. $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$. Because optimization with Adam is faster, the number of epochs for SVHN and CIFAR-10 is reduced to 20 and 15, respectively. While the performance gap is smaller with this optimizer, it is still significant and suggests that our methods are beneficial with other optimizers than the simple SGD. Note that this analysis is fairly superficial. The thorough investigation of the interplay between aggregators and momentum-based optimizers is a compelling future research direction.
(a) SVHN: Training loss
(b) SVHN: Update similarity to the SGD update
(c) CIFAR-10: Training loss
(d) CIFAR-10: Update similarity to the SGD update
Figure 10: Results with the Adam optimizer.
Appendix E Computation time
In this appendix, we compare the computation time of SGD with that of SSJD for all the aggregators that we experimented with. Since we used the same architecture for MNIST, Fashion-MNIST and Kuzushiji-MNIST, we only report the results for one of them. It is important to clarify all of the factors that affect this computation time. First, the batch size affects the number of rows in the Jacobian to aggregate. Increasing the batch size thus requires more GPU memory and the aggregation of a taller matrix. Then, some aggregators, e.g. $\mathcal{A}_{\text{Nash-MTL}}$ and $\mathcal{A}_{\text{MGDA}}$, seem to greatly increase the run time. When the aggregation is the bottleneck, a faster implementation will be necessary to make them usable in practice. Lastly, the current implementation of our library for JD with PyTorch is still fairly inefficient in terms of memory management, which in turn limit how well the GPU can parallelize. Therefore, these results just give a rough indication of the current computation times, but they are in no way final.
Table 7: Time required in seconds for one epoch of training with SGD and different instances of SSJD, on an NVIDIA L4 GPU. The batch size is always 32.
Method SVHN CIFAR-10 EuroSAT MNIST
SGD 0.79 0.50 0.81 0.47
SSJD + $\mathcal{A}_{\text{Mean}}$ 1.41 1.76 2.93 1.64
SSJD + $\mathcal{A}_{\text{MGDA}}$ 5.50 5.22 6.91 5.22
SSJD + $\mathcal{A}_{\text{DualProj}}$ 1.51 1.88 3.02 1.76
SSJD + $\mathcal{A}_{\text{PCGrad}}$ 2.78 3.13 4.18 3.01
SSJD + $\mathcal{A}_{\text{GradDrop}}$ 1.57 1.90 3.06 1.78
SSJD + $\mathcal{A}_{\text{IMTL-G}}$ 1.48 1.79 2.94 1.69
SSJD + $\mathcal{A}_{\text{CAGrad}}$ 1.93 2.26 3.42 2.17
SSJD + $\mathcal{A}_{\text{RGW}}$ 1.42 1.76 2.89 1.73
SSJD + $\mathcal{A}_{\text{Nash-MTL}}$ 7.88 8.12 9.33 7.91
SSJD + $\mathcal{A}_{\text{Aligned-MTL}}$ 1.53 1.98 2.97 1.71
SSJD + $\mathcal{A}_{\text{UPGrad}}$ 1.80 2.01 3.21 1.90
Appendix F Forward propagation of the Gramian through some typical layers
In this appendix, we study the forward propagation of the Gramian, as per (12), through some typical neural network layers. We start by noting that some computations can be saved when the function has a block structure.
Let ${\boldsymbol{g}}:{\mathbb{R}}^{n}\times{\mathbb{R}}^{\ell}\to{\mathbb{R}}^{m}$ a function that maps an input ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$ and some parameter ${\boldsymbol{p}}\in{\mathbb{R}}^{\ell}$ to some output ${\boldsymbol{g}}({\boldsymbol{x}},{\boldsymbol{p}})$. Applying ${\boldsymbol{g}}$ independently in parallel yields a function ${\boldsymbol{f}}:{\mathbb{R}}^{n\times k}\times{\mathbb{R}}^{\ell}\to{\mathbb{R}}^{m\times k}$, where $k$ is the number of blocks. [^6] The block size $k$ could be, among others, the batch size or the number of elements of an activation. For an input $X\in{\mathbb{R}}^{n\times k}$ and a parameter ${\boldsymbol{p}}\in{\mathbb{R}}^{\ell}$,
${\boldsymbol{f}}(X,{\boldsymbol{p}})$ =[||$\ {\boldsymbol{g}}(X\cdot{\boldsymbol{e}}_{1},{\boldsymbol{p}})&\dots&{\boldsymbol{g}}(X\cdot{\boldsymbol{e}}_{k},{\boldsymbol{p}})\\ |&&|\end{bmatrix}.$
In that case, forwarding the Gramian requires the matrices
$\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}(X,{\boldsymbol{p}})$ $=\begin{bmatrix}\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{g}}(X\cdot{\boldsymbol{e}}_{1},{\boldsymbol{p}})&0&\cdots&0\\ 0&\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{g}}(X\cdot{\boldsymbol{e}}_{2},{\boldsymbol{p}})&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{g}}(X\cdot{\boldsymbol{e}}_{k},{\boldsymbol{p}})\end{bmatrix}$
$\mathcal{G}\kern-1.0pt_{2}{\boldsymbol{f}}(X,{\boldsymbol{p}})$ $=\begin{bmatrix}J_{1}^{\vphantom{\top}}J_{1}^{\top}&J_{1}^{\vphantom{\top}}J_{2}^{\top}&\cdots&J_{1}^{\vphantom{\top}}J_{k}^{\top}\\ J_{2}^{\vphantom{\top}}J_{1}^{\top}&J_{2}^{\vphantom{\top}}J_{2}^{\top}&\cdots&J_{2}^{\vphantom{\top}}J_{k}^{\top}\\ \vdots&\vdots&\ddots&\vdots\\ J_{k}^{\vphantom{\top}}J_{1}^{\top}&J_{k}^{\vphantom{\top}}J_{2}^{\top}&\cdots&J_{k}^{\vphantom{\top}}J_{k}^{\top}\\ \end{bmatrix},$
where $J_{i}^{\vphantom{\top}}=\mathcal{J}\kern-2.0pt_{2}{\boldsymbol{g}}(X\cdot{\boldsymbol{e}}_{i},{\boldsymbol{p}})$.
Because of the block-diagonal structure of $\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}(X,{\boldsymbol{p}})$, for a given $G\in\mathcal{M}_{nk}$, the term $\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}(X,{\boldsymbol{p}})~{}\cdot~{}G\cdot~{}\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}(X,{\boldsymbol{p}})^{\top}$ can be computed in $\mathcal{O}(k^{2}(mn^{2}~{}+~{}nm^{2}))$ operations rather than the naive $\mathcal{O}(k^{3}(mn^{2}~{}+~{}nm^{2}))$. Appendices F.1, F.2 and F.3 provide considerations about Gramian forwarding through some usual layers that have a particular block structure. For parameterized layers, $\mathcal{G}\kern-1.0pt_{2}f(X,{\boldsymbol{p}})$ can sometimes be precomputed efficiently, depending on the layer. More details are given in Appendix F.4 for a fully connected layer.
F.1 Activation functions
Activation functions generally consider each element of the activation independently. We thus have $m=n=1$, $\ell=0$ and $k$ equal to the number of elements of the activation multiplied by the batch size. The matrix $\mathcal{J}\kern-2.0pt_{1}f(X)$ is thus diagonal. The computational complexity of forwarding a Gramian through such a function is then $\mathcal{O}(k^{2})$. This applies to most usual activation functions, such as ReLU, ELU, sigmoid, and tanh.
F.2 Softmax layer
In the softmax layer, the group size $k$ is the batch size. Moreover, $m=n$ and $\ell=0$. The computational complexity of forwarding a Gramian is thus $\mathcal{O}(k^{2}m^{3})$.
F.3 BatchNorm layer
The batch normalization layer [57], in its simplest form (without the learnable affine transformation), normalizes its input on the batch dimension. Here $k$ is the input size, $\ell=0$, and $m=n$ is the batch size. In that case, the computational complexity of forwarding a Gramian is $\mathcal{O}(k^{2}m^{3})$.
F.4 Fully connected layer
In a fully connected layer, ${\boldsymbol{g}}:{\mathbb{R}}^{n}\times\mathbb{R}^{m\times n}\times{\mathbb{R}}^{m}\to{\mathbb{R}}^{m}$ is given by ${\boldsymbol{g}}({\boldsymbol{x}},W,{\boldsymbol{b}})=W{\boldsymbol{x}}+{\boldsymbol{b}}$ and ${\boldsymbol{f}}$ has a $k$-blockwise structure with $k$ the batch size. If $X\in{\mathbb{R}}^{n\times k}$ and $G\in{\mathbb{R}}^{nk\times nk}$, then forwarding those through the layer yields
$\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}(X,W,{\boldsymbol{b}})\cdot G\cdot\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}(X,W,{\boldsymbol{b}})^{\top}+\mathcal{G}\kern-1.0pt_{2}{\boldsymbol{f}}(X,W,{\boldsymbol{b}})+\mathcal{G}\kern-1.0pt_{3}{\boldsymbol{f}}(X,W,{\boldsymbol{b}})$
with
$\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}(X,W,{\boldsymbol{b}})$ $=\begin{bmatrix}W&0&\cdots&0\\ 0&W&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&W\end{bmatrix}$
$=I_{k}\otimes W,$
where $\otimes$ is the Kronecker product and $I_{k}$ the identity matrix of dimension $k$. In principle, $\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}(X,W,{\boldsymbol{b}})\cdot G\cdot\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}(X,W,{\boldsymbol{b}})^{\top}$ can be computed efficiently.
The Gramians with respect to the parameters are also rather simple. Let ${\boldsymbol{x}}_{i}=X{\boldsymbol{e}}_{i}$, then
$\mathcal{G}\kern-1.0pt_{2}{\boldsymbol{f}}(X,W,{\boldsymbol{b}})$ $=\begin{bmatrix}{\boldsymbol{x}}_{1}^{\top}{\boldsymbol{x}}_{1}I_{m}&{\boldsymbol{x}}_{1}^{\top}{\boldsymbol{x}}_{2}I_{m}&\cdots&{\boldsymbol{x}}_{1}^{\top}{\boldsymbol{x}}_{k}I_{m}\\ {\boldsymbol{x}}_{2}^{\top}{\boldsymbol{x}}_{1}I_{m}&{\boldsymbol{x}}_{2}^{\top}{\boldsymbol{x}}_{2}I_{m}&\cdots&{\boldsymbol{x}}_{2}^{\top}{\boldsymbol{x}}_{k}I_{m}\\ \vdots&\vdots&\ddots&\vdots\\ {\boldsymbol{x}}_{k}^{\top}{\boldsymbol{x}}_{1}I_{m}&{\boldsymbol{x}}_{k}^{\top}{\boldsymbol{x}}_{2}I_{m}&\cdots&{\boldsymbol{x}}_{k}^{\top}{\boldsymbol{x}}_{k}I_{m}\\ \end{bmatrix}$
$=(X^{\top}X)\otimes I_{m}$
$\mathcal{G}\kern-1.0pt_{3}{\boldsymbol{f}}(X,W,{\boldsymbol{b}})$ $=\begin{bmatrix}I_{m}&I_{m}&\cdots&I_{m}\\ I_{m}&I_{m}&\cdots&I_{m}\\ \vdots&\vdots&\ddots&\vdots\\ I_{m}&I_{m}&\cdots&I_{m}\\ \end{bmatrix}$
$=1_{k}\otimes I_{m},$
with $1_{k}$ the all $1$ matrix of dimension $k\times k$.
Therefore, (12) rewrites as
$(I_{k}\otimes W)\cdot G\cdot(I_{k}\otimes W)+(X^{\top}X+1_{k})\otimes I_{m}.$
This illustrates that one should think of $G$ as the quadratic mapping $J\to JGJ^{\top}$, and similarly of $J$ as the linear mapping $G\to JGJ^{\top}$. Representing those mappings with arrays would be sub-optimal. Taking advantage of their structure (in this case, the Kronecker product of two matrices) would be beneficial.
It is worth noting that even though $(X^{\top}X+1_{k})\otimes I_{m}$ encompasses the dependencies of the Gramian to the parameters, it does not require any computation over the parameter dimension. If the number of parameters is large, this can provide a substantial improvement over the naive computation of the Gramian via the Jacobian.